### v1

# HIP-LLM: A Hierarchical Imprecise Probability Approach to Reliability Assessment of Large Language Models

Abstract
Large Language Models (LLMs) are increasingly deployed across diverse domains, raising the need for rigorous reliability assessment methods. Existing benchmark-based evaluations primarily offer descriptive statistics of model accuracy over datasets, providing limited insight into the probabilistic behavior of LLMs under real operational conditions. This paper introduces HIP-LLM, a Hierarchical Imprecise Probability framework for modeling and inferring LLM reliability. Building upon the foundations of software reliability engineering, HIP-LLM defines LLM reliability as the probability of failure-free operation over a specified number of future tasks under a given Operational Profile (OP). HIP-LLM represents dependencies across (sub-)domains hierarchically, enabling multi-level inference from subdomain to system-level reliability. HIP-LLM embeds imprecise priors to capture epistemic uncertainty and incorporates OPs to reflect usage contexts. It derives posterior reliability envelopes that quantify uncertainty across priors and data. Experiments on multiple benchmark datasets demonstrate that HIP-LLM offers a more accurate and standardized reliability characterization than existing benchmark and state-of-the-art approaches. A publicly accessible repository of HIP-LLM is provided.
keywords:
Large Language Model; Reliability; Hierarchical Bayesian Inference; Operational Profile; Epistemic Uncertainty; Imprecise Probability
journal: Reliability Engineering & System Safety
[…]
1 Introduction
Large language models (LLMs) are increasingly applied across a wide range of tasks, from general-purpose reasoning to highly specialized applications. For instance, recent studies [1, 2, 3, 4, 5, 6] employ LLMs as tools to support reliability analysis and safety assurance for safety-critical systems. This expanding scope of application underscores the urgent need for comprehensive and rigorous evaluation methods to assess the reliability of the LLMs themselves, given the high-stakes nature of those applications. While numerous studies have proposed domain-specific performance evaluations in areas such as coding [7, 8, 9], medicine [10], and system safety [11, 12, 13], others have focused on broader properties of LLMs, including safety (avoiding harmful content) [14, 15, 16], robustness (resisting adversarial attacks such as jailbreaks) [17, 18], fairness [19], and privacy [20, 21]. However, there remains a notable lack of research dedicated to reliability, which motivates this work.
While software reliability is highly interrelated with the aforementioned properties (a formal definition and its distinction to other properties will be discussed in later sections), it is a unique property that emphasizes the failure probabilities, specified operational time and conditions, and statistically valid uncertainty quantification. As a “user-centric” property [22, 23, 24], the delivered reliability depends on the behavior of end-users and how the software is expected to be used in practice. This notion is made measurable through the Operational Profile (OP) [25, 26], which specifies a probability distribution over the types of demands that users actually place on the software. The recent report from OpenAI [27] provide the most comprehensive data on how user-groups use ChatGPT, which aligns with the idea of OPs.
Despite software reliability assessment has been studied for decades [28, 29], state-of-the-art LLM evaluation largely relies on benchmarks [30]. This creates several research gaps:
  • Gap-1
    (static benchmarks vs dynamic operational use): While we acknowledge the usefulness of benchmarks, they are primarily designed for purposes such as comparing and ranking LLMs on given tasks under fixed datasets, neglecting the variability and dynamic nature of user’s behaviors in real-world applications. Thus, benchmarks cannot serve the purpose of assessing reliability, especially when the OP of the users is different from the data distribution (implicitly) represented by the benchmark dataset.
  • Gap-2
    (independent benchmarks vs hierarchical dependencies): While LLMs are generally designed for broad use (covering various task types such as legal reasoning and coding), these tasks naturally exist at different levels of abstraction. It is often useful to organize them hierarchically, dividing the input space into domains and subdomains [31]. High-level domains [^1] While the precise definition of what constitutes a domain or subdomain is an application-oriented question that may vary case by case reflecting the assessor’s domain knowledge, in this reliability modeling work we do not prescribe specific taxonomies. Instead, we introduce a general hierarchical modeling structure that can accommodate different dependency assumptions (cf. Section 5 for more discussion). (e.g., law vs. coding) can be treated as largely independent, whereas subdomains within a domain (e.g., coding in Python vs. coding in C++) are more likely to exhibit dependencies. In such cases, performance in one subdomain may influence users’ confidence in another, whereas cross-domain effects are minimal. Reliability assessment should therefore reflect these dependency relationships and support evaluation at multiple abstraction levels, depending on the assessor’s role. For instance, a LLM vendor may be interested in system-wide reliability across all domains, a programmer may focus on reliability in coding tasks, and a Python developer may be specifically concerned with the reliability of Python coding. Existing benchmarks, however, typically evaluate LLM capabilities either in isolation or by simply averaging performance across tasks, that obscure these dependencies relationships [31].
  • Gap-3
    (descriptive statistics vs statistical inference): While descriptive statistics summarize the observed dataset by reporting point-estimate benchmark scores, they do not formally model how the data were generated. Even when variance is reported as a way of “quantifying uncertainty”, it quantifies variability within the sample itself. In contrast, statistical inference treats the observed results as samples from an underlying stochastic process or population and uses probabilistic modeling to estimate the parameters of that process [32]. This inferential perspective enables generalization beyond the fixed benchmark dataset and supports principled uncertainty quantification about reliability which concerns a underlying population of future inputs [31].
Indeed, acknowledging these research gaps, recent studies [32, 31] attempt to bridge them by applying both “frequentist” and Bayesian statistical inference methods. However, despite representing important first steps, they do not fully resolve the aforementioned gaps (e.g., not explicitly incorporate the OP) and introduce new gaps:
  • Gap-4
    (failure probability vs future failure-free runs) While both [31, 32] examined the metric of failure probability [^2] We later show that this is a special case of our reliability model., they did not extend it to the more practical and widely adopted definition of reliability that predicts the probability of failure-free runs over a specified future operational time [33, 34, 35].
  • Gap-5
    (non-informative vs informative priors): The work [32] is a frequentist approach that cannot embed prior knowledge explicitly like Bayesian methods. The work [31] employs non-informative priors in its Bayesian models, without facilitating the embedding of any informative prior knowledge that end-users may have.
To bridge the five research gaps collectively, we propose HIP-LLM, a hierarchical imprecise probability approach to assess LLM reliability. HIP-LLM devises a more rigorous and versatile assessment with the following key features. First, similar to [31], it structures the model into independent high-level domains, with each domain further divided into dependent subdomains. In this way, reliability can be evaluated at multiple levels of detail while preserving statistical dependencies, so that information from one subdomain contributes to the inference about other subdomains within the same domain. We then incorporate OPs as operational weights at each level of the hierarchy (subdomain $\rightarrow$ domain $\rightarrow$ general-purpose LLM) to reflect the dynamic operational use of subdomains and domains in practice. In contrast to [31], which relies on non-informative priors, HIP-LLM enables the incorporation of informative prior knowledge. Since selecting a single prior in Bayesian inference is often controversial due to its encoding of epistemic uncertainty, we adopt Imprecise Probability [36, 37] that represents uncertainty about the prior itself without committing to a single distribution. Consequently, reliability metrics (defined in terms of both failure probability and the probability of failure-free runs in future operations) are expressed through posterior distribution envelopes at different levels of the hierarchy, reflecting both uncertainties from the data and the prior knowledge.
In summary the contributions of this work are as follows:
  • 1.
    We formally define the reliability assessment problem for LLMs in accordance with established software reliability standards, while delineating its distinction from related properties.
  • 2.
    We propose HIP-LLM, a hierarchical imprecise probability model that explicitly addresses key research gaps in the current state-of-the-art, by modeling OPs, hierarchical dependencies, statistical principled uncertainty quantification on reliability (both failure probability and future failure-free runs over specified time), and imprecise prior knowledge.
  • 3.
    We release a public repository containing all experimental data, source code, and models at https://github.com/aghazadehchakherlou-web/llm-imprecise-bayes.
The remainder of the paper is organized as follows. Section 2 introduces preliminaries and reviews related works on LLM assessment. Section 3 provides the modeling details of HIP-LLM. Section 4 evaluates and illustrates HIP-LLM via experiments on datasets. Section 5 discusses the limitations and assumptions, and finally Section 6 concludes the paper.
2 Preliminaries and Related Works
2.1 Software Reliability
According to American National Standards Institute (ANSI), software reliability is defined as [38]:
Definition 1 (Software reliability).
The probability of failure-free software operation for a specified period of time in a specified environment
This probabilistic definition is also largely adopted by software reliability engineering literature [28], which makes reliability amenable to statistical modeling and permits risk-aware aggregation across heterogeneous operational uses. The specific probabilistic metric used to model reliability is domain-dependent and determined by the operational nature of the software. For example, continuous-time systems are continuously operated in the active control of a process, whereas on-demand systems are only invoked upon receipt of discrete demands. In the latter case, such as nuclear power protection systems, the probability of failure on demand (pfd) has been adopted in standards [23, 39, 40] and extensively studied [41, 33, 42].
Although related, reliability is distinct from safety, security, and accuracy. Reliability focuses on the probability of failures (relative to a specification), a stochastic and usage-weighted concept by OPs [42]. Safety concerns critical failures with catastrophic consequences, i.e., whether a failure can lead to unacceptable harm. For instance, a system can be reliable but unsafe (if frequent but non-harmful failures are tolerated while rare failures produce catastrophic harm), or safe but unreliable (if it fails frequently but never produce catastrophic harm by fail-safe design) [^3] We note that for systems in which all failures are catastrophic (e.g., safety-critical systems), safety and reliability assessments do not require different statistical reasoning, such as when measuring the probability that a system will operate safely (i.e., without critical failures) over a given mission.. Security is concerned with resisting malicious threats: it normally assumes an explicit threat model and seeks to prevent or mitigate intentional compromises). While security fundamentally requires the analysis of a malicious threat actor, reliability and safety can be studied within benign operational environments where failure originates from sources such as design flaws and implementation bugs. Within the AI/ML community, accuracy is typically a narrower performance metric, defined as a descriptive statistic (proportion of correct predictions) over a fixed and given dataset. However, such a dataset does not necessarily represent the unknown ground-truth population of inputs that the model will face upon deployment, and accuracy does not quantify the uncertainties of failure-free operations over future time like reliability.
A central insight from software reliability engineering is that reliability is user-context dependent: the same software can exhibit very different delivered reliability under different patterns of use. The OP formalizes this dependency, which is defined as [23, 25, 28]:
Definition 2 (Operational Profile (OP)).
An OP is a probability distribution of inputs to a software system, representing the relative frequencies with which different input sequences expected occur during actual operation.
There are some caveats regarding the term “inputs” in the definition of OP, particularly when considering factors such as software memory and dependencies [^4] In practice, because software systems may possess memory, their success/failure often depends on entire sequences of inputs rather than isolated ones. Accordingly, “inputs” may denote a complete stimulus, which may encompass the full sequence of interactions within a task, or even the cumulative sequence of inputs since the system was last reinitialized.. We refer readers to [23] for a detailed discussion.
In the context of LLMs, OPs serve an analogous role [27]: they specify the distribution of task types (e.g., factual question answering, coding, summarization) and their relative importance in deployment. Incorporating OPs into LLM reliability analysis ensures that evaluation metrics capture not just aggregate benchmark scores, but also how well a model performs across the realistic mix of tasks it is expected to face.
2.2 Benchmark-Based Evaluation of LLMs via Descriptive Statistics
Most evaluation methods for LLMs report only point estimates such as accuracy scores. Question-answering benchmarks such as MMLU [43] and RACE [44] measure how often a model picks the right answer in multiple-choice questions. Code benchmarks like DS-1000 [8], InterCode [7], and HumanEval Pro/MBPP Pro [9] test whether generated programs run correctly. Assistant and autonomy benchmarks such as GAIA [45] and H-CAST [46] look at overall task completion—whether the system can finish a complex job successfully. Model cards and reports for Claude 3 [47], GPT-4o [48], and code models [49] publish these benchmark results as single numbers per model. Broader frameworks like HELM [50] add other measures such as calibration, robustness, fairness, and efficiency, but still mainly give point values. In real-world use, Clio [21] collects statistics on what kinds of tasks people actually do with AI with a focus on privacy. Method studies like multilingual chain-of-thought [51] and ReAct [52] suggest ways to improve reasoning and interaction, judged by benchmark accuracy.
Recent studies widely acknowledge that benchmark-based evaluations for LLMs are valuable but inherently limited due to the new characteristics and challenges introduced by LLMs. Benchmarks facilitate progress tracking and model comparison, yet they capture only a partial view of LLM capabilities. Chang et al. [30] note that traditional benchmarks (reusing fixed task sets for ranking models) have driven early advances but are now approaching their limits due to data repetition, task saturation [^5] Over time, models are trained on similar data, inflating scores and diminishing benchmark usefulness., and limited assessment of reasoning or interactivity. Liu et al. [53] emphasize that standard benchmarks often overlook moral, safety, and social dimensions of model behavior. In domain-specific contexts, Croxford et al. [10] show that medical benchmarks miss crucial aspects such as factual accuracy and clinical reasoning, highlighting the need for domain prior knowledge. Saleh et al. [54] observe that efficiency and performance benchmarks remain fragmented and inconsistent. From a safety perspective, Liu et al. [55] warn that toxicity and robustness tests are often too narrow and easily gamed, recommending flexible, real-world evaluations. Finally, Ye et al. [56] introduce psychometric principles, arguing that current benchmarks frequently lack checks for reliability and fairness.
While we acknowledge the usefulness of LLM benchmarks for progress tracking of LLM versions, comparison and ranking of LLMs, they can be used to make claims on LLM reliability. Our HIP-LLM approach goes beyond benchmarks, by framing the reliability assessment of LLMs as statistical inference problems.
2.3 Statistical Inference for Evaluating LLMs
Benchmark-based descriptive statistics simply summarize the collected data, whereas statistical inference draws probabilistic conclusions about the underlying population or process that generated those data, thereby enabling generalization, prediction and uncertainty quantification. In the AI/ML community, many studies have proposed statistically principled approaches to quantify uncertainty for individual inputs, such as text prompts for LLMs, using methods like conformal prediction [57, 58]. However, these approaches focus on instance-level uncertainty and do not capture uncertainty at the operational reliability level. To the best of our knowledge, only two existing works extend uncertainty quantification to this broader reliability perspective for LLMs.
In the Anthropic report [32], Miller frames LLM evaluations as statistical experiments and argues for reporting uncertainty alongside benchmark scores. Using classical inference tools, the study constructs confidence intervals via the Central Limit Theorem, applies clustered standard errors for correlated items, recommends paired inference for model comparisons, and develops power analysis and variance reduction techniques. This work emphasizes that benchmark outcomes should be treated as data drawn from an underlying population, moving beyond descriptive statistics.
HiBayES [31] applies hierarchical Bayesian generalized linear models that: (a) use a Bayesian statistical inference (multilevel Binomial/Poisson Generalized Linear Models) and captures the nested structure of LLM evaluations; (b) employ partial pooling to account for dependencies across levels; (c) provide a fully probabilistic framework, yielding full posterior distributions to quantify uncertainties, rather than relying on point-estimates.
Despite these advances, important limitations remain. Neither Miller’s frequentist framework nor HiBayES explicitly defines/models the OP (even though both implicitly assume that the data represent it). Moreover, neither approach can effectively incorporate (potentially imprecise) prior knowledge: Miller’s work is restricted to classical frequentist inference, while HiBayES relies on non-informative priors. Both also focus on the narrower notion of failure probability as reliability, overlooking the more general and standardized reliability definition in terms of the probability of failure-free runs over future operations. The latter requires predictive inference that accounts for the propagation of uncertainties into the future: a more demanding but practical reliability claim [33, 34].
2.4 Robust Bayesian analysis
Robust Bayesian analysis represents a general framework for investigating the sensitivity of posterior measures to uncertainties in the inputs of Bayesian inference [59, 60]. Regarding uncertainties in priors, while several dedicated methods have been proposed [61, 33, 62, 63], Imprecise Probability has emerged as one of the most widely adopted approaches [36, 37, 64, 65, 66]. It addresses the problem of prior uncertainty by avoiding reliance on a single prior distribution, rather representing a credal set of plausible priors and deriving posterior bounds that reflect this epistemic uncertainty.
To illustrate the main idea of the Imprecise Probability framework, consider a simple coin-flipping problem where we wish to estimate the probability of heads $\theta$ with data $D$ ($n=10$, $k=3$ heads). As shown in Table 1, instead of a single point estimate or a single Beta posterior distribution, the posterior envelope (i.e., a set of posterior distributions) of Imprecise Probability [^6] For simplicity and illustrative purpose, the example used here is based on the model of imprecise Linearly Updated Conjugate prior Knowledge (iLUCK) [65]) which leverages the conjugacy for analytical posterior results. Our HIP-LLM is not using this iLUCK model due to the non-linearity of the hierarchical probabilistic model proposed in Section 3. reflects both the observed data and the epistemic uncertainty arising from imprecise prior knowledge (represented by the set of Beta priors).
Table 1: Classical Bayesian vs Imprecise Probability for a coin-flipping example.
Feature Classical Bayesian Imprecise Probability
Prior $\mathrm{Beta}(\alpha=2,\beta=2)$ $\mathrm{Beta}(\alpha,\beta),\;\alpha\in[1,3],\beta\in[1,3]$
Posterior $\theta\mid D\sim\mathrm{Beta}(5,9)$ $\theta\mid D\sim\mathrm{Beta}(3+\alpha,7+\beta),\;\alpha\in[1,3],\beta\in[1,3]$
Poster. mean $\mathbb{E}[\theta\mid D]=0.36$ $\mathbb{E}[\theta\mid D]\in[0.31,0.38]$
Our hierarchical solution, HIP-LLM, is a robust Bayesian approach that addresses the aforementioned gaps. Similar to HiBayES, it adopts a hierarchical structure, but instead of yielding a single posterior, it reports posterior envelopes over imprecise priors and uncertain OPs (as variables).
3 The Model: HIP-LLM
The proposed method HIP-LLM stands for Hierarchical Imprecise Probability for Large Language Models reliability assessment. Before introducing the framework, we first formally define the reliability of LLMs.
LLMs are software; therefore, to ensure compatibility with the more general and standardized definition of software reliability (Def. 1), we define LLM reliability as follows:
Definition 3 (LLM reliability).
The probability that an LLM produces failure-free responses over a specified number of future tasks (sequences of closely related queries), under specified (sub-)domains and operational environment.
The definition of LLM reliability closely parallels the standardized definition of software reliability, as both emphasize probabilistic reasoning and the requirement of failure-free performance in future operations. The key distinctions arise from the nature of LLMs as “on-demand” software. Whereas classical software reliability is typically framed in terms of “operational time” that covers continuously operated systems like controllers over clock-time, LLM reliability is defined more explicitly in terms of a specified number of future discrete tasks (each consisting of sequences of related queries). Furthermore, while both definitions account for operational conditions, the LLM context requires explicit reference to domains and subdomains, reflecting the general-purpose and multi-domain design of LLMs. Finally, in traditional software, failure is typically defined as a deviation from the specification, whereas there is no specification for LLMs and thus “failures of LLMs” [^7] A complete and formal characterization of what constitutes a “failure” for LLMs remains an open research question, and out of the scope of this paper. Cf. Section 5 for discussions. are often informally/implicitly characterized by factual errors (hallucinations) [67, 68] or divergence from human-expert answers [69, 70].
3.1 Problem Formulation
While we provide a table of notations in C, according to Def. 3, LLM reliability can be formalised as follows:
Definition 4 (Formalised LLM reliability).
Let $\mathcal{X}$ denote the input-space of all possible tasks for a given LLM, and let $\pi$ be the OP, i.e., a probability distribution over $\mathcal{X}$ that reflects the likelihood of encountering each task $x\in\mathcal{X}$ in practice. Consider a sequence of $n\geq 1$ tasks independently and identically distributed (i.i.d.) according to the OP $\pi$, let $\mathbb{I}(x_{t})\in\{0,1\}$ indicate success (1) or failure (0) on the $t$-th task $x_{t}$, then the LLM reliability is:
$R(n,\pi)=\Pr_{x_{t}\sim\pi}\left(\bigcap_{t=1}^{n}\{\mathbb{I}(x_{t})=1\}\right)$ (1)
Intuitively, $R(n,\pi)$ represents the probability that the LLM will operate failure-free across the next $n$ i.i.d. tasks according to the OP $\pi$. Importantly, we make the following notes:
  1. 1.
    A key assumption of the above definition is the task failures/success are i.i.d. Bernoulli trials. Such modeling is not uncommon in reliability modeling, especially for critical on-demand systems. A typical justification is when demands are rare, and the states/memory of the software and its operational environment are effectively “reset” in-between [23, 71]. In the context of LLMs, we carefully define our reliability metric in terms of i.i.d “tasks”, rather than individual prompts which are often contextually dependent. A task may consist of a sequence of related prompts aimed at achieving a single task goal. We acknowledge that modern LLMs (e.g., ChatGPT) typically retain chat history as contextual memory, which can violate the independence assumption between tasks. However, most LLMs also provide the option to start a new chat session (for a new single task) without any memory and history from previous chat sessions, aligning with the assumption of “resetting” the LLM.
  2. 2.
    While $R(n,\pi)$ is the general form of future reliability of processing $n$ tasks, the special case $1-R(1,\pi)$ (where $n=1$) represents the failure probability [33, 34, 35] that studied by, e.g., [31].
  3. 3.
    Since the input-space $\mathcal{X}$ represents all possible LLM tasks, so $R(n,\pi)$ is the general-purpose reliability of the LLM under study. For (sub-)domain specific reliability, we need to partition $\mathcal{X}$ and derive “local” OPs; then the lower level (sub-)domain specific reliability can be similarly derived like Eq. (1).
To rigorously assess the formally defined LLM reliability, coping with the aforementioned notes, the next subsection introduces our proposed solution HIP-LLM. It models the LLM as a hierarchical structure consisting of independent domains, each containing statistically dependent subdomains (cf. Fig. 1). Here the modeled dependencies and independencies represent the epistemic structure of our hierarchical Bayesian model. That is, observing failures in one (sub-)domain may or may not update our beliefs about the reliability of other (sub-)domains. One possible example of justification [^8] Cf. Section 5 for discussions on the validity of this hierarchal dependency assumption. is: We model coding and law as independent because they rely on distinct competencies of the LLM—coding on formal, symbolic reasoning and syntax manipulation, and law on narrative understanding and normative interpretation. Since these skills draw from largely separate representations and training data, failures in one domain provide little information about failures in the other. In Bayesian terms, their failure probabilities can be treated as a priori independent parameters, reflecting separate latent skill dimensions of the model. On the other hand, sub-domains failure probabilities are modeled as dependent parameters given their shared LLM competencies.
Our goal is to infer the posterior distributions of future reliability (and its special case the failure probability), at the subdomain, domain, and general-purpose LLM levels, based on observed correct responses from tasks within each subdomain.
Figure 1: Schematic representation of the hierarchical LLM, domain, and subdomain structure for reliability estimation fo M LLM models.
3.2 Proposed Solution
Consider a hierarchical structure of an LLM comprising independent domains $D_{1},D_{2},\dots,$ $D_{m}$, where each domain $D_{i}$ contains statistically dependent subdomains $S_{i1},\dots,S_{in_{i}}$ (Fig. 2). We wish to infer the posterior distributions over subdomain, domain and LLM level reliabilities by observing $C_{ij}$ correct responses out of $N_{ij}$ trials (tasks) in each subdomain.
Fig. 2 presents a detailed view of the hierarchical structure (subdomains $\rightarrow$ domains $\rightarrow$ LLM), with assumed priors and parameters. We assume that subdomain reliabilities within a domain are dependent through a shared prior, and that domain reliabilities are aggregated from their subdomains according to task-specific operational profiles. Our goal is to construct a principled hierarchical Bayesian model that supports information sharing across dependent subdomains through partial pooling (Sec. 3.2.1) and uncertainty quantification via Imprecise Probability (Sec. 3.2.2).
Figures 1 and 2 illustrate a general hierarchical structure comprising multiple LLM instances ($LLM^{(1)},LLM^{(2)},\ldots,LLM^{(M)}$). For clarity, however, Theorems 16 focus on the reliability assessment of a single LLM system. Accordingly, we omit the superscript $(k)$ and use unindexed symbols ($D_{i}$, $S_{ij}$, $p_{L}$). Extending the framework to multiple LLMs is straightforward—apply it to each system independently and compare their posterior distributions.
Figure 2: Hierarchical structure with independent domains and dependent subdomains. For readability, parameters and priors are shown only for one subdomain $S_{ij}$ and its parent domain $D_{i}$ under a representative $LLM_{k}$; the remaining subdomains, domains ($i=1,\dots,m$), and LLMs $(k=1,\dots,M)$ are identical, differing only in their indices.
3.2.1 Hierarchical Bayesian Framework for LLM Reliability Modeling
As seen in Fig. 2, inference proceeds hierarchically from the bottom up. At the subdomain level, the observed data consist of the number of correct responses $C_{ij}$ out of total trials $N_{ij}$. These update the subdomain reliabilities $\theta_{ij}$, which are modeled with a Binomial likelihood [^9] The assumption of independent Bernoulli trials with constant success probability $\theta_{ij}$ may not capture all real dependencies, but it serves as an effective approximation for modeling subdomain outcomes [32, 31, 62]. Aggregating across $N_{ij}$ trials then yields the Binomial likelihood, whose support $\{0,\dots,N_{ij}\}$ matches the possible counts of correct responses observed in subdomain $S_{ij}$. $C_{ij}\mid\theta_{ij},N_{ij}\;\sim\;\mathrm{Binomial}(N_{ij},\theta_{ij}).$ and a Beta prior [^10] We place a Beta prior on subdomain reliability: $\theta_{ij}\sim\mathrm{Beta}(\alpha_{i},\beta_{i}).$ The Beta distribution is the most common choice for probabilities bounded in $[0,1]$, and it is conjugate to the Binomial likelihood, ensuring closed-form updates and computational stability.. The Beta prior is parameterized by domain-level hyperparameters $(\mu_{i},\nu_{i})$ [^11] To make priors more intuitive and interpretable, we use a reparameterization. Instead of specifying the Beta prior directly in terms of $(\alpha_{i},\beta_{i})$, we express it as $\theta_{ij}\mid\mu_{i},\nu_{i}\;\sim\;\mathrm{Beta}(\mu_{i}\nu_{i},(1-\mu_{i})\nu_{i}),$ where $\mu_{i}$ denotes the expected reliability (prior mean, $\mu_{i}=\mathbb{E}[\theta_{ij}\mid\mu_{i},\nu_{i}]$) and $\nu_{i}$ denotes the prior strength, reflecting the confidence in $\mu_{i}$ (equivalent sample size or pseudo-counts, $\nu_{i}=\alpha_{i}+\beta_{i}$). This reparameterization makes prior beliefs easier to specify and justify., representing the expected reliability and the prior strength within domain $D_{i}$. These hyperparameters are in turn governed by domain-specific hyperpriors: $\mu_{i}\sim\mathrm{Beta}(a_{i},b_{i})$ [^12] This choice reflects the idea that $\mu_{i}$ itself is a probability lying in $(0,1)$, and the Beta distribution provides a flexible family of shapes that can express different prior beliefs about domain reliability, ranging from diffuse to highly concentrated around particular values (e.g., favoring higher values, lower values, or balanced around 0.5)., and $\nu_{i}\sim\mathrm{Gamma}(c_{i},\text{rate}=d_{i})$ [^13] This treats $\nu_{i}$ as a positive random variable ($\nu_{i}>0$) reflecting how tightly subdomains within a domain are assumed to cluster around $\mu_{i}$. When $\nu_{i}$ is small, the prior is diffuse and allows substantial variation across subdomains (weak pooling). When $\nu_{i}$ is large, the prior concentrates mass near $\mu_{i}$ and subdomains are tightly clustered (strong pooling). [^14] The Gamma distribution is commonly written in two equivalent forms: shape–rate and shape–scale. We use the shape–rate form, $\nu_{i}\sim\mathrm{Gamma}(c_{i},\mathrm{rate}=d_{i})$, with $\mathbb{E}[\nu_{i}]=c_{i}/d_{i}$ and $\mathrm{Var}[\nu_{i}]=c_{i}/d_{i}^{2}$. If a library expects the shape–scale form, set $\theta_{i}=1/d_{i}$ and write $\nu_{i}\sim\mathrm{Gamma}(c_{i},\mathrm{scale}=\theta_{i})$; the two parameterizations are mathematically identical under $\theta_{i}=1/d_{i}$. The only practical concern is clarity and reproducibility—accidentally treating a rate as a scale (or vice versa) would change the prior’s mean and variance..
Because the hyperparameters $(\mu_{i},\nu_{i})$ are shared across all subdomains in domain $D_{i}$, the information is pooled: subdomains with few observations are stabilized by drawing on evidence from other subdomains in the same domain, while subdomains with many observations are influenced mainly by their own data. This hierarchical setup achieves the desired dependence within domains, while domains remain independent (cf. Section 5 for discussions on this setup).
Once posterior subdomain reliabilities $\theta_{ij}$ are inferred, they are aggregated into domain-level reliabilities $p_{i}=\sum_{j}\Omega_{ij}\theta_{ij}$ using OP weights $\Omega_{ij}$, which reflect the practical importance of subdomains. These domain-level posteriors are then further combined into the overall LLM reliability, $p_{L}=\sum_{i}W_{i}p_{i}$, using domain operational weights $W_{i}$. Note, OPs at different levels are represented by the variables $\Omega_{ij}$ and $W_{i}$, which, for a given LLM use case, can be instantiated either as fixed constants (when usage is certain) or as probability distributions that encode uncertainty about how the LLM will be used in practice [72, 73, 74, 75].
3.2.2 Uncertainty Handling via Imprecise Probability
To address the epistemic uncertainty in prior specification, we adopt an Imprecise Probability approach: we specify the hyper-hyper-parameters $a_{i},b_{i},c_{i},d_{i}$ as intervals (Eq. 3.2.2) rather than point values. This produces posterior envelopes (lower and upper bounds) at the subdomain, domain, and overall LLM levels.
$a_{i}\in[a_{i}^{\min},a_{i}^{\max}],\quad b_{i}\in[b_{i}^{\min},b_{i}^{\max}],$
$c_{i}\in[c_{i}^{\min},c_{i}^{\max}],\quad d_{i}\in[d_{i}^{\min},d_{i}^{\max}]$ (2)
At the subdomain level, we compute posterior bounds for each task type by considering all admissible hyperparameter configurations:
$\underline{Pr}(\theta_{ij}\mid C_{i})\leq Pr(\theta_{ij}\mid C_{i})\leq\overline{Pr}(\theta_{ij}\mid C_{i})$ (3)
Note, $C_{i}$ denotes the data is domain $i$. Since we consider dependencies among subdomains, the posterior $\theta_{ij}$ is a function of all data $C_{i}$ in domain $i$.
Later (Theorem. 2 and Theorem. 3) we will discuss that the closed-form densities (like subdomain level) do not exist for domain and LLM level reliability. We represent the cumulative distribution function (CDF) envelopes for the domain and overall LLM levels.
At the domain level, uncertainty propagates upward through OP weights $p_{i}=\sum_{j}\Omega_{ij}\theta_{ij}$, producing domain reliability bounds:
$\underline{F}_{p_{i}}(t\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{p_{i}}(t\mid C_{i},h_{i}),\quad\overline{F}_{p_{i}}(t\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{p_{i}}(t\mid C_{i},h_{i}),\quad t\in[0,1]$
The domain level posterior $p_{i}$ only considers $C_{i}$, given our assumption on cross-domain independence.
At the LLM level, uncertainty aggregates across all $k$ domains through domain weights $p_{L}=\sum_{i=1}^{k}W_{i}p_{i}$, while respecting cross-domain independence, resulting in system-level reliability bounds:
$\underline{F}_{p_{L}}(t\mid\text{data})=\inf_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{p_{L}}(t\mid\text{data},\mathcal{H}),$
$\overline{F}_{p_{L}}(t\mid\text{data})=\sup_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{p_{L}}(t\mid\text{data},\mathcal{H})$
where, $t\in[0,1]$.
While in the next subsection, we develop Theorems 1–3 of deriving posterior sets for those non-failure probability variables $\theta_{ij}$s, $p_{i}$s and $p_{L}$, the posterior distribution sets for future reliability of passing next $n_{F}$ tasks at each level, e.g., for a domain $i$:
$\underline{Pr}((\sum_{j}\Omega_{ij}\theta_{ij})^{n_{F}}\mid C_{i})\leq Pr((\sum_{j}\Omega_{ij}\theta_{ij})^{n_{F}}\mid C_{i})\leq\overline{Pr}((\sum_{j}\Omega_{ij}\theta_{ij})^{n_{F}}\mid C_{i})$ (4)
can also be derived, as shown in our Theorems 4–6.
3.2.3 Theorems
The following theorems are the main mathematical results of HIP-LLM. Intuitively, given the imprecise prior knowledge encoded by the hyperparameters, the theorems derive the posterior distributions of non-failure probabilities at different hierarchical levels, as well as the future reliability, based on the probabilistic reasoning model illustrated in Fig. 2, conditioned on the observed task failure data.
Theorem 1 (Sub-domain level non-failure probability).
For subdomain $S_{ij}$ in domain $D_{i}$, let $C_{i}=\{(C_{ik},N_{ik})\}_{k=1}^{n_{i}}$ be the observed data. Let the admissible set of hyperparameters be
$\mathcal{A}_{i}\;=\;[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}],$
and write $h_{i}=(a_{i},b_{i},c_{i},d_{i})$. Then, for any $h_{i}\in\mathcal{A}_{i}$, the marginal posterior density of $\theta_{ij}$ is
$Pr(\theta_{ij}\mid C_{i},h_{i})=\frac{f_{\mathrm{marg}}(\theta_{ij},C_{i};h_{i})}{Z_{\mathrm{marg}}(h_{i})},$
where $f_{\mathrm{marg}}$ (unnormalized posterior) and $Z_{\mathrm{marg}}$ (normalizing constant) are
$\hskip 28.45274ptf_{\mathrm{marg}}(\theta_{ij},C_{i};h_{i})=$
$\hskip 56.9055pt\int_{0}^{1}\int_{0}^{\infty}\Bigg[\prod_{k\neq j}\int_{0}^{1}d\theta_{ik}\Bigg]L(\boldsymbol{\theta}_{i})\,Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i})\,Pr(\mu_{i},\nu_{i}\mid h_{i})\,d\mu_{i}\,d\nu_{i}$
$Z_{\mathrm{marg}}(h_{i})=\int_{0}^{1}\int_{0}^{\infty}Pr(C_{i}\mid\mu_{i},\nu_{i})\,Pr(\mu_{i},\nu_{i}\mid h_{i})\,d\mu_{i}\,d\nu_{i},$
with $L(\boldsymbol{\theta}_{i})=Pr(C_{i}\mid\boldsymbol{\theta}_{i})$.
The imprecise marginal posterior is characterized by the lower/upper envelopes
$\underline{Pr}(\theta_{ij}\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}Pr(\theta_{ij}\mid C_{i},h_{i}),\qquad\overline{Pr}(\theta_{ij}\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}Pr(\theta_{ij}\mid C_{i},h_{i}).$
The proof of Theorem 1 is presented at Appendix A.2. In this theorem, the subdomain posterior $Pr(\theta_{ij}\mid C_{i},h_{i})$ has a closed-form density because of conjugacy: the Beta prior combined with the Binomial likelihood yields a mixture of Beta distributions after marginalizing over the hyperparameters $(\mu_{i},\nu_{i})$, which can be expressed and evaluated as a proper probability density function.
However, for later Theorems closed-form densities do not exist because $p_{i}=\sum_{j}\Omega_{ij}\theta_{ij}$ and $p_{L}=\sum_{i}W_{i}p_{i}$ are weighted sums of dependent random variables—the distribution of a sum of Beta random variables has no analytical form except in trivial cases [^15] Assuming the weights are constants. When the weights are modeled as random variables with their own probability distributions, the same problem persists (if not harder).. Computing such densities would require intractable multi-dimensional integrals. The Cumulative Distribution Function (CDF) formulation sidesteps this problem: $F_{p_{i}}(t\mid C_{i},h_{i})=\int_{0}^{1}\int_{0}^{\infty}F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}$ only requires a two-dimensional integral over $(\mu_{i},\nu_{i})$, where the conditional CDF can be computed via Monte Carlo sampling of independent Betas. Since CDFs provide all necessary information for practical reliability assessment (probabilities, quantiles, expectations), they are the natural representation when densities are unavailable.
Theorem 2 (Domain level posterior non-failure probability).
For domain $D_{i}$ with local OP weights $\Omega_{ij}$ (where $\sum_{j=1}^{n_{i}}\Omega_{ij}=1$), let $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ be the domain-level non-failure probability. Define the admissible set of hyper-hyper-parameters
$\mathcal{A}_{i}=[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}],$
and write $h_{i}=(a_{i},b_{i},c_{i},d_{i})$.
Then, for any $h_{i}\in\mathcal{A}_{i}$, the posterior distribution of $p_{i}$ is characterized by its cumulative distribution function:
$F_{p_{i}}(t\mid C_{i},h_{i})$ $=\Pr(p_{i}\leq t\mid C_{i},h_{i})$
$\hskip 56.9055pt=\int_{0}^{1}\int_{0}^{\infty}F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}$
where
  • 1.
    $F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})$ is the conditional CDF of $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ given that $\theta_{ij}\mid\mu_{i},\nu_{i},C_{i}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Beta}(C_{ij}+\mu_{i}\nu_{i},N_{ij}-C_{ij}+(1-\mu_{i})\nu_{i})$ for $j=1,\ldots,n_{i}$,
  • 2.
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})$ is the hyper-posterior obtained via Bayes’ rule:
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})=$
    $\hskip 28.45274pt\frac{Pr(C_{i}\mid\mu_{i},\nu_{i})\,\text{Beta}(\mu_{i}\mid a_{i},b_{i})\,\text{Gamma}(\nu_{i}\mid c_{i},\text{rate}=d_{i})}{\int_{0}^{1}\int_{0}^{\infty}Pr(C_{i}\mid\mu,\nu)\,\text{Beta}(\mu\mid a_{i},b_{i})\,\text{Gamma}(\nu\mid c_{i},\text{rate}=d_{i})\,d\mu\,d\nu}$
The imprecise domain posterior is characterized by CDF envelopes:
$\underline{F}_{p_{i}}(t\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{p_{i}}(t\mid C_{i},h_{i}),\qquad\overline{F}_{p_{i}}(t\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{p_{i}}(t\mid C_{i},h_{i}).$
For the proof details, we refer reads to A.3.
Theorem 3 (LLM-level posterior non-failure probability).
For the LLM system with domain weights $W_{i}$ (where $\sum_{i=1}^{m}W_{i}=1$), let $p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$ be the LLM-level failure probability and $\text{data}=\{C_{1},\ldots,C_{m}\}$ the observed data across all domains. Assume cross-domain independence.
Define the domain-level admissible sets
$\mathcal{A}_{i}=[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}]$
and write $h_{i}=(a_{i},b_{i},c_{i},d_{i})$ for $i=1,\ldots,m$. Define the LLM-level admissible set as the Cartesian product
$\mathcal{A}_{\text{LLM}}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{m},$
and collect the domain hyperparameters as $\mathcal{H}=(h_{1},\ldots,h_{m})\in\mathcal{A}_{\text{LLM}}$.
Then, for any $\mathcal{H}\in\mathcal{A}_{\text{LLM}}$, the posterior distribution of $p_{L}$ is characterized by its cumulative distribution function:
$F_{p_{L}}(t\mid\text{data},\mathcal{H})=Pr(p_{L}\leq t\mid\text{data},\mathcal{H})=$
$\hskip 42.67912pt\int\cdots\int G\big(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data}\big)\prod_{i=1}^{m}Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\prod_{i=1}^{m}d\mu_{i}\,d\nu_{i}$
where
  • 1.
    $G(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data})$ is the conditional CDF of $p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$ given all hyperparameters, defined as
    $G\big(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data}\big)=\int_{\mathcal{R}_{L}(t)}\prod_{i=1}^{m}f_{p_{i}}(p_{i}\mid\mu_{i},\nu_{i},C_{i})\,dp_{1}\cdots dp_{m},$
    where $\mathcal{R}_{L}(t):=\{(p_{1},\ldots,p_{m})\in(0,1)^{m}:\sum_{i=1}^{m}W_{i}p_{i}\leq t\}$, and $f_{p_{i}}(\cdot\mid\mu_{i},\nu_{i},C_{i})$ is the conditional density of $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ under $\theta_{ij}\mid\mu_{i},\nu_{i},C_{i}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Beta}(C_{ij}+\mu_{i}\nu_{i},N_{ij}-C_{ij}+(1-\mu_{i})\nu_{i})$,
  • 2.
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})$ is the domain-level hyper-posterior for domain $i$:
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})=$
    $\hskip 28.45274pt\frac{Pr(C_{i}\mid\mu_{i},\nu_{i})\,\text{Beta}(\mu_{i}\mid a_{i},b_{i})\,\text{Gamma}(\nu_{i}\mid c_{i},\text{rate}=d_{i})}{\int_{0}^{1}\int_{0}^{\infty}Pr(C_{i}\mid\mu,\nu)\,\text{Beta}(\mu\mid a_{i},b_{i})\,\text{Gamma}(\nu\mid c_{i},\text{rate}=d_{i})\,d\mu\,d\nu}$
  • 3.
    Cross-domain independence ensures $Pr(\{\mu_{i},\nu_{i}\}_{i=1}^{m}\mid\text{data},\mathcal{H})=\prod_{i=1}^{m}Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})$.
The imprecise LLM posterior is characterized by CDF envelopes:
$\underline{F}_{p_{L}}(t\mid\text{data})=\inf_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{p_{L}}(t\mid\text{data},\mathcal{H})$
$\overline{F}_{p_{L}}(t\mid\text{data})=\sup_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{p_{L}}(t\mid\text{data},\mathcal{H})$
For more details, see Section A.4.
Theorems 1, 2, and 3 characterize the posterior distribution of non-failure probability $\theta_{ij}$ or aggregated $p_{i}$, $p_{L}$. However, in practice, we often care about the reliability over a specified number of consecutive future operations, e.g., “what is the probability that an LLM succeeds on the next 10 tasks in a row?” or “what is the probability that the LLM succeeds on the next 20 coding tasks?”.
The following set of theorems extends the aforementioned theorems to characterize the full posterior distribution of reliability for $n_{F}$ consecutive future operations, i.e., the probability that the LLM operates $n_{F}$ consecutive failure-free tasks. Conditioning is on the observed evaluation data across subdomains $S_{ij}$: $C_{ij}$ correct generations out of $N_{ij}$ prompts.
Theorem 4 (Subdomain posterior reliability for $n_{F}$ future operations).
For subdomain $S_{ij}$ in domain $D_{i}$, let $C_{i}=\{(C_{ik},N_{ik})\}_{k=1}^{n_{i}}$ denote all observed data in the domain. Let the admissible set of hyperparameters be
$\mathcal{A}_{i}=[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}],$
and write $h_{i}=(a_{i},b_{i},c_{i},d_{i})$. Define the reliability random variable [^16] Comparing to the reliability definition in Def. 4, the OPs are omitted as we assume they are fixed constants in these theorems. $R_{ij}(n_{F})=\theta_{ij}^{n_{F}}$, representing the probability of $n_{F}$ consecutive failure-free operations in subdomain $S_{ij}$.
For any $h_{i}\in\mathcal{A}_{i}$, the posterior cumulative distribution function of $R_{ij}(n_{F})$ is:
$F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i})=\Pr(\theta_{ij}^{n_{F}}\leq r\mid C_{i},h_{i})=\int_{0}^{r^{1/n_{F}}}Pr(\theta_{ij}\mid C_{i},h_{i})\,d\theta_{ij},$
where $Pr(\theta_{ij}\mid C_{i},h_{i})$ is the marginal posterior density from Theorem 1.
The imprecise posterior distribution is characterized by the CDF envelopes:
$\underline{F}_{R_{ij}(n_{F})}(r\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i}),$
$\overline{F}_{R_{ij}(n_{F})}(r\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i})$
The transformation $\theta_{ij}\mapsto\theta_{ij}^{n_{F}}$ generally does not yield a closed-form density. We therefore characterize $R_{ij}(n_{F})$ through its CDF. See A.5 computational methods.
Theorem 5 (Domain posterior reliability for $n_{F}$ future operations).
For domain $D_{i}$ with local OP weights $\Omega_{ij}$ (where $\sum_{j=1}^{n_{i}}\Omega_{ij}=1$), let $C_{i}=\{(C_{ik},N_{ik})\}_{k=1}^{n_{i}}$ denote all observed data in the domain. Let the admissible set of hyperparameters be
$\mathcal{A}_{i}=[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}],$
and write $h_{i}=(a_{i},b_{i},c_{i},d_{i})$. Define the domain-level reliability
$p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij},\quad R_{i}(n_{F})=p_{i}^{n_{F}},$
representing the probability of $n_{F}$ consecutive failure-free operations at the domain level.
For any $h_{i}\in\mathcal{A}_{i}$, the posterior distribution of $R_{i}(n_{F})$ is characterized by its CDF:
$F_{R_{i}(n_{F})}(r\mid C_{i},h_{i})=\Pr\left(\left[\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}\right]^{n_{F}}\leq r\,\bigg|\,C_{i},h_{i}\right),$
computed by integrating over the joint posterior $Pr(\boldsymbol{\theta}_{i}\mid C_{i},h_{i})$.
The imprecise posterior distribution is characterized by the CDF envelopes:
$\underline{F}_{R_{i}(n_{F})}(r\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{R_{i}(n_{F})}(r\mid C_{i},h_{i}),$
$\overline{F}_{R_{i}(n_{F})}(r\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{R_{i}(n_{F})}(r\mid C_{i},h_{i})$
The CDF can be computed via numerical integration over the joint posterior $Pr(\boldsymbol{\theta}_{i}\mid C_{i},h_{i})$ and Monte Carlo sampling, cf. A.5.
Theorem 6 (LLM posterior reliability for $n_{F}$ future operations).
For the LLM system with an OP of domain weights $W_{i}$ (where $\sum_{i=1}^{m}W_{i}=1$), let $\text{data}=\{C_{1},\ldots,C_{m}\}$ denote all observed data across domains. Assume cross-domain independence.
Define the domain-level admissible sets
$\mathcal{A}_{i}=[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}]$
and write $h_{i}=(a_{i},b_{i},c_{i},d_{i})$ for $i=1,\ldots,m$. Define the LLM-level admissible set as the Cartesian product
$\mathcal{A}_{\text{LLM}}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{m},$
and collect the domain hyperparameters as $\mathcal{H}=(h_{1},\ldots,h_{m})\in\mathcal{A}_{\text{LLM}}$.
Define the LLM-level reliability
$p_{L}=\sum_{i=1}^{m}W_{i}p_{i},\quad R_{L}(n_{F})=p_{L}^{n_{F}},$
representing the probability of $n_{F}$ consecutive failure-free operations at the LLM level.
For any $\mathcal{H}\in\mathcal{A}_{\text{LLM}}$, the posterior distribution of $R_{L}(n_{F})$ is characterized by its CDF:
$F_{R_{L}(n_{F})}(r\mid\text{data},\mathcal{H})=\Pr\left(\left[\sum_{i=1}^{m}W_{i}p_{i}\right]^{n_{F}}\leq r\,\bigg|\,\text{data},\mathcal{H}\right),$
computed by integrating over $\prod_{i=1}^{m}Pr(p_{i}\mid C_{i},h_{i})$.
The imprecise posterior distribution is characterized by the CDF envelopes:
$\underline{F}_{R_{L}(n_{F})}(r\mid\text{data})=\inf_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{R_{L}(n_{F})}(r\mid\text{data},\mathcal{H}),$
$\overline{F}_{R_{L}(n_{F})}(r\mid\text{data})=\sup_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{R_{L}(n_{F})}(r\mid\text{data},\mathcal{H})$
Again, while the closed-form density is not available, the CDF can be computed via numerical integration and Monte Carlo sampling. See A.5 for details.
4 Evaluation
To demonstrate and evaluate our HIP-LLM, we empirically investigate five research questions (RQs) in this section.
4.1 Research Questions
RQ1 (Effectiveness): How effectively can HIP-LLM assess and compare posterior reliability distributions across different levels of the hierarchy, considering uncertainties propagated from subdomains to domains, and finally to general-purpose LLMs? In this RQ, we aim to demonstrate the use case of our HIP-LLM as a reliability assessment tool.
RQ2 (Sensitivity to hyperparameters): How sensitive are the posterior estimates of HIP-LLM to the hyperparameters ($a_{i},b_{i},c_{i},d_{i}$)? These hyperparameters represent the assessors’ (imprecise) prior knowledge, thus understanding their sensitivity to the posteriors may provide insights on how prior knowledge can be elicited.
RQ3 (Sensitivity to OPs): How sensitive are the posterior reliability estimates of HIP-LLM to variations in the OPs that characterize operational usage of LLMs at various levels? The delivered and perceived reliability of an LLM depends on how it will be used, i.e. the OP. General-purpose LLMs may exhibit lower sensitivity to variations in OPs, whereas LLMs trained for specific (sub-)domains are likely to be more sensitive. We investigate and demonstrate how HIP-LLM can characterize and quantify such OP-dependent reliability variations.
RQ4 (Predictability): How can HIP-LLM predicts future reliability of passing next $n_{F}$ tasks? While RQ1–RQ3 focus on failure probabilities, which is a special case of reliability, we additionally aim to demonstrate HIP-LLM’s capability to predict the probability of successfully completing $n_{F}$ future tasks and to quantify how this reliability varies as the reliability requirement $n_{F}$ changes.
4.2 Experimental Setup
OPs
Since this work represents an initial academic study on reliability assessment modeling for LLMs without actual operational field data, we follow a similar approach to [31, 32] by leveraging existing benchmark datasets. To emphasize the role of OPs, we “simulate” the OPs by assigning probability distributions over these datasets for sampling and by specifying operational weights across (sub-)domains. Without loss of generality, and consistent with our three-level hierarchy, we define the task-level OP as a uniform distribution and assign operational weights at the subdomain and domain levels proportionally to their dataset sizes. That said, assessors may use alternative distributions as the 3-level OPs when additional information is available, such as those approximated from historical usage data [24, 76] or user behavior reports [27].
Data
Specifically, we evaluate our hierarchical framework by simulating LLM operational data from four widely used benchmarks, structured into two domains with two subdomains each (same as HiBayEs [31]), instantiated by the following datasets:
  • 1.
    Domain 1 (Coding)
    • (a)
      $\text{Subdomain}_{11}$: basic Python programming tasks with unit-test based evaluation (MBPP [9]).
    • (b)
      $\text{Subdomain}_{12}$: data-science oriented Python problems involving libraries such as pandas and numpy (DS-1000 [8]).
  • 2.
    Domain 2 (Reasoning):
    • (a)
      $\text{Subdomain}_{21}$: reading comprehension problems where the LLM must answer yes/no questions given short passages (BoolQ [77]).
    • (b)
      $\text{Subdomain}_{22}$: high school level reading comprehension problems with multiple-choice answers (RACE-H [44]).
All experiments were conducted using publicly available APIs of GPT-4o, GPT-4o-mini (from OpenAI), Claude Sonnet 4.5, and Claude Haiku 3.5 (from Anthropic) [^17] For academic research purposes only which is permitted under both vendors’ terms of service for research publications.. Table 2 reports sample accuracies per subdomain, i.e., the proportions of correct responses under Pass@1181818Produce 1 solution per task. Score as correct if that solution passes verification.. To operationalize the i.i.d. Bernoulli trail assumption on tasks, all LLM evaluations were conducted with cleared context between tasks. Each task in MBPP, DS-1000, BoolQ, and RACE-H was processed as an independent API call with no conversational history, ensuring that outcomes are not influenced by previous interactions.
Table 2: Evaluation results using Pass@1. Entries are per-subdomain accuracies ($C_{ij}/N_{ij}$). Rightmost column shows the row mean across models.
Dom Subdom (Data.)
4o-mini
4o
sonnet-4.5
haiku-3.5
Mean
$\text{Dom}_{1}$ $\text{Subdom}_{11}$ (MBPP) 0.440 0.471 0.450 0.447 0.452
$\text{Subdom}_{12}$ (DS-1000) 0.490 0.420 0.493 0.483 0.472
Mean 0.465 0.446 0.472 0.465
$\text{Dom}_{2}$ $\text{Subdom}_{21}$ (BoolQ) 0.890 0.909 0.900 0.883 0.896
$\text{Subdom}_{22}$ (RACE-H) 0.820 0.552 0.840 0.859 0.768
Mean 0.855 0.731 0.87 0.871
LLM Mean 0.661 0.585 0.671 0.668
From the benchmark accuracy scores, all LLMs appear to perform better in the Reasoning domain than in the Coding domain. Their performances across most subdomains are similar, except for RACE-H, where model OpenAI-4o shows a clear weakness. The small variance between accuracy values for corresponding subdomains across models suggests that the performance differences are more domain-driven than model-driven.
Hyperparameters
All the hyperparameters we are using for generating the set of figures in the next subsection is shown in B.
4.3 Results and Analysis
This section reports and analyses the empirical results to the RQs.
4.3.1 RQ1 (Effectiveness)
Fig. 3 shows posterior CDF envelopes for the four subdomains (MBPP, DS–1000, BoolQ, RACE–H) across the four LLMs where a right–shifted CDF indicates higher reliability and a tighter band indicates greater certainty.
Coding tasks
On MBPP (top left), the envelopes for GPT–4o–mini and Haiku 3.5 are nearly indistinguishable and lie to the left of GPT–4o, while Sonnet 4.5 is right–most but still overlapping with–4o. On DS–1000 (top right), GPT–4o–mini lies clearly to the left, while Haiku 3.5 overlaps almost entirely with GPT–4o—their envelopes coincide so closely that Haiku 3.5 is barely distinguishable. Sonnet 4.5 remains right–most, with only partial overlap with GPT–4o and Haiku 3.5.
Reasoning tasks
On BoolQ (bottom left), GPT–4o–mini is almost identical to Haiku 3.5 on the left, while Sonnet 4.5 and GPT–4o nearly coincide on the right. Hence the envelopes separate into two close pairs: $\text{mini}\approx\text{Haiku 3.5}<\text{Sonnet 4.5}\approx 4\text{o}$. On RACE–H (bottom right), Haiku 3.5 lies clearly left–most, GPT–4o–mini overlaps partly with Haiku but extends rightward, followed by GPT–4o, and finally Sonnet 4.5 on the far right.
(a) Subdomain11
(b) Subdomain12
(c) Subdomain21
(d) Subdomain22
Figure 3: Posterior CDF envelopes at subdomain level. Rows correspond to domains (D1 (coding): MBPP, DS-1000; D2 (reasoning): BoolQ, RACE-H).
Fig. 4 aggregates subdomains within each domain via the operational weights ($\Omega_{ij}$) and reports the posterior CDFs of the domain reliabilities ($p_{i}$). Note, for simplicity, we just assign the operational weights ($\Omega_{ij}$) proportionally according to the dataset sizes of the two sub-domains in each domain $i$. Similarly to sub-domain results, we may observe and compare domain-level non-failure probabilities $p_{i}$. Again, comparing to accuracy scores (point estimates) and HiBayEs (single posterior distributions), our HIP-LLM yields distribution envelopes that considers more types of uncertainties.
(a) Domain1
(b) Domain2
Figure 4: Posterior CDF envelopes for domain-level reliabilities $p_{i}=\sum_{j}\Omega_{ij}\theta_{ij}$. Operational profile weights are $\Omega_{1\cdot}=(0.204,0.796)$ for D1 and $\Omega_{2\cdot}=(0.483,0.517)$ for D2.
Fig. 5 aggregates both domains into the overall LLM reliability $p_{L}$ using the cross–domain operational weights W=[0.149, 0.851]W=[0.149,\,0.851] for [Coding, Reasoning]. As before, the weights are simply assigned proportionally according to the dataset sizes of domains. The figure shows, 4o-mini and Haiku bands overlap almost completely, indicating near–equivalent reliability, while GPT–4o and Sonnet 4.5 partially overlap, reflecting moderate but consistent uncertainty between them. Overall, Sonnet 4.5 remains most reliable (which is consistent with Table 2).
Figure 5: Posterior CDF envelope for overall LLM reliability $p_{L}=\sum_{i}W_{i}p_{i}$ with domain weights W=(0.149, 0.851)W=(0.149,\,0.851).
4.3.2 RQ2 (Sensitivity to hyperparameters)
(a) Sensitivity of LLM posterior reliability distribution to $a_{i}$.
(b) Sensitivity of LLM posterior reliability distribution to $b_{i}$.
(c) Sensitivity of LLM posterior reliability distribution to $c_{i}$.
(d) Sensitivity of LLM posterior reliability distribution to $d_{i}$.
Figure 6: Effect of variations in the hyperparameter values ($a$, $b$, $c$, and $d$) on the overall LLM reliability.
[…]
Fig. 6(a) illustrates the sensitivity of the LLM-level posterior reliability distributions to the hyperparameter $a_{i}$, which serves as the first shape parameter of the Beta prior $\mu_{i}\sim\mathrm{Beta}(a_{i},b_{i})$. Increasing $a_{i}$ effectively shifts the prior mean of $\mu_{i}=a_{i}/(a_{i}+b_{i})$ upward, representing stronger prior belief in higher domain reliability. The rightward shift indicates a higher central estimate of LLM-level reliability.
Fig. 6(b) presents the corresponding sensitivity analysis for the hyperparameter $b_{i}$, the second shape parameter of the same Beta prior on $\mu_{i}$. Increasing $b_{i}$ decreases the prior mean $\mu_{i}=a_{i}/(a_{i}+b_{i})$, thereby encoding a more pessimistic prior expectation of domain reliability. The leftward displacement indicates a reduction in the posterior mean reliability.
Fig. 6(c) and Fig. 6(d) together illustrate how the LLM-level posterior reliability distribution responds to changes in the hyperparameters $c_{i}$ and $d_{i}$, which jointly define the Gamma prior $\nu_{i}\sim\mathrm{Gamma}(c_{i},\text{rate}=d_{i})$ controlling the pooling strength among subdomains within a domain. Increasing $c_{i}$ raises the expected prior strength $E[\nu_{i}]=c_{i}/d_{i}$ and reduces its relative variance, enforcing stronger coupling across subdomains so that their reliabilities $\theta_{ij}$ concentrate more closely around the domain mean $\mu_{i}$; this yields a modest rightward shift in the cumulative distribution, indicating a more stable and slightly higher central reliability estimate. Conversely, increasing $d_{i}$ lowers $E[\nu_{i}]$ and weakens pooling, permitting greater variability among subdomains and consequently broadening and shifting the posterior envelopes leftward. This broader, left-shifted distribution reflects increased epistemic uncertainty and the greater influence of weaker subdomains on the aggregate, leading to more conservative and less certain reliability assessments.
4.3.3 RQ3 (Sensitivity to OPs)
(a) Anthropic Sonnet 4.5
(b) OpenAI 4-o
Figure 7: Posterior CDF envelopes for Reasoning-domain ($D_{2}$) reliability under alternative operational weights on RACE–H, $\Omega_{\mathrm{RACE}}\in\{0.10,\,0.517,\,0.90\}$ with $\Omega_{\mathrm{BoolQ}}=1-\Omega_{\mathrm{RACE}}$. (a) Sonnet 4.5; (b) OpenAI 4-o.
Fig. 7 shows posterior CDF envelopes for the Reasoning-domain reliability $p_{2}$ under three operational profiles, $\Omega_{\mathrm{RACE}}\in\{0.10,\,0.517,\,0.90\}$ with $\Omega_{\mathrm{BoolQ}}=1-\Omega_{\mathrm{RACE}}$. Fig. 7(a) reports Anthropic Sonnet 4.5; Fig. 7(b) reports OpenAI 4-o.
(a) Sonnet 4.5. As $\Omega_{\mathrm{RACE}}$ increases the CDF shifts slightly left (lower delivered $p_{2}$), with strong overlap among bands. This weak sensitivity to the operational profile is consistent with the subdomain accuracies being both high and similar ($\theta_{\text{BoolQ}}\approx 0.90$ vs. $\theta_{\text{RACE-H}}\approx 0.84$): the mixture $p_{2}=\Omega_{\mathrm{BoolQ}}\theta_{\text{BoolQ}}+\Omega_{\mathrm{RACE}}\theta_{\text{RACE-H}}$ changes only modestly as weight moves from BoolQ to RACE–H. The mild widening of the envelope at higher $\Omega_{\mathrm{RACE}}$ reflects slightly larger posterior uncertainty on RACE–H relative to BoolQ.
(b) OpenAI 4-o. In contrast, the envelopes are well separated and move substantially left as $\Omega_{\mathrm{RACE}}$ increases. This strong dependence on the operational weighting arises because 4-o’s subdomain accuracies differ widely ($\theta_{\text{BoolQ}}\approx 0.91$ vs. $\theta_{\text{RACE-H}}\approx 0.55$), so the mixture $p_{2}$ is dominated by whichever subdomain receives more weight.
Overal, the posterior CDF envelopes for OpenAI 4-o exhibit a strong dependence on the BoolQ/RACE–H weighting, whereas Sonnet 4.5 is comparatively invariant to $\Omega_{\mathrm{RACE}}$. This arises because GPT-4-o’s subdomain accuracies differ widely ($0.91$ vs $0.55$), producing a hierarchical posterior with pronounced weight sensitivity. These results suggest that GPT-4-o’s reasoning performance is more uneven across benchmark types, while Sonnet-4.5 demonstrates domain-level robustness.
4.3.4 RQ4 (Predictability)
Fig. 8(a) presents the expected LLM-level reliability $\mathbb{E}[R_{L}(n_{F})]$ as a function of the operational horizon $n_{F}$, where $n_{F}$ denotes the number of consecutive operations in the future. The horizontal axis displays $n_{F}$ on a logarithmic scale, while the vertical axis shows the expected reliability—the posterior mean probability that the LLM successfully completes all next $n_{F}$ consecutive tasks without failure.
(a) Expected LLM reliability $\mathbb{E}[R_{L}(n_{F})]$ versus operational horizon $n_{F}$. Imprecise posterior envelopes show decay in reliability over consecutive operations of four models.
(b) Envelope widths (upper–lower) quantify epistemic uncertainty, peaking at intermediate horizons and shrinking as reliability saturates.
Figure 8: Reliability and uncertainty of four LLMs across operational horizons.
At $n_{F}=1$ all models exhibit high expected reliability, $\mathbb{E}[R_{L}(1)]\approx 0.80\text{--}0.87$, i.e., a high probability of producing a correct answer on the very next task (the non‐failure probability). As the horizon length $n_{F}$ increases, expected reliability drops sharply for every model. Across all horizons the ordering is stable: Sonnet-4.5 achieves the highest reliability, GPT-4o follows closely, while Haiku-3.5 and GPT-4o-mini form the lower pair (with GPT-4o-mini typically the lowest). However, as shown in As Fig. 8(b), the uncertainty (measured by the interval width) of Sonnet-4.5’s reliability prediction is also the highest.
As Fig. 8(b) shows, at $n_{F}=1$ the envelopes are quite tight, indicating low uncertainty in one–step reliability. The bands widen modestly for small–to–medium horizons (i.e., when $N_{F}$ increases from 2 to $\sim 6$), as a result of compounded prediction uncertainties. As $n_{F}$ increases, all curves converge toward zero, indicating reduced uncertainty in the reliability prediction. Intuitively, this suggests greater confidence that the LLMs are unlikely to pass all $n_{F}$ future tasks.
5 Discussion
In this section, we discuss the assumptions and limitations of the proposed HIP-LLM.
Assumptions on hierarchical dependencies of failure probabilities
Inspired by HiBayES [31], we introduce a hierarchical dependency structure where dependencies are defined at the level of failure probabilities across related task groups. Specifically, high-level domains (e.g., coding vs. law) are modeled as independent because their underlying competencies and training data sources are substantially distinct. In contrast, subdomains within a domain (e.g., coding in C++ vs. coding in Python) are modeled as having dependent failure probabilities, reflecting shared latent skills, representations, or reasoning patterns that influence performance across related tasks. This abstraction allows tractable modeling of correlated reliability while acknowledging that the true dependencies among knowledge areas are more complex and fluid. While we do not prescribe specific taxonomies of domain/sub-domains for LLMs in this paper, our formulation is intended to be without loss of generality: users of HIP-LLM can redefine their own “domains” to reflect their own assumptions about dependence. For example, if users believe that two domains (say, A and B) are not fully independent, they may treat them as two subdomains under a newly introduced higher-level synthetic domain. Such flexibility allows HIP-LLM to adapt to different operational interpretations or empirical evidence of dependency, reflecting the assessor’s domain knowledge.
Note, such (in)dependencies at the parameter level (i.e., correlations or independence between the failure probabilities of domains or subdomains) are different from task-level outcome dependencies, which we discuss next.
Assumptions of i.i.d. Bernoulli Trails for Task Outcomes
It is important to distinguish the assumption of independent task outcomes from the (in)dependencies among failure probabilities discussed earlier. The latter concern parameter-level relationships, i.e., how the long-run failure rates across domains or subdomains may co-vary due to shared competencies, whereas the former concerns instance-level independence of individual task results, given those parameters. Even if the failure probabilities of two sub-domains are correlated random variables, we can still assume that, conditional on those probabilities, each task outcome within a sub-domain is an independent Bernoulli trial.
As discussed earlier in Def. 4, the i.i.d. assumption is pragmatically justified in the context of LLM evaluations when tasks are executed in isolation, such as initiating new chat sessions without contextual memory or cross-task influence. Nevertheless, caution is warranted: in practice, the way tasks are designed, sequenced, or batched can introduce subtle dependencies that violate this idealization. Assessors should therefore ensure that testing procedures (e.g., resetting chat sessions) are aligned with the independence assumption.
Indeed, the i.i.d. Bernoulli trial assumption remains a simplification, motivated by mathematical convenience and interpretability. Future work should explicitly incorporate uncertainty or “doubt” about this assumption to examine how sensitive reliability claims are to potential task-level dependencies of outcomes [71].
What constitutes a failure for LLMs
Reliability, by definition, concerns “failures”. In traditional software systems, failures are relatively straightforward to define, as they can be explicitly linked to violations of formal specifications or requirements. However, for AI systems, and particularly for LLMs, no such explicit specification exists. As a result, a complete and formal characterization of what constitutes a failure for LLMs remains an open research challenge. The notion of “failure” can vary across applications and stakeholders. For instance, factual errors, inconsistency [^19] Most existing LLM evaluation studies focus on one-shot accuracy metrics (e.g., Pass@1), whereas real-world usage often involves iterative prompting. Inconsistencies across multiple outputs can undermine user trust and thus regarded as a failure., harmful or biased content, or divergence from human judgment may all be considered failures depending on the context [30].
That said, in this work, our focus is not on prescribing a specific definition of LLM failure but on providing a reliability assessment modeling framework that is agnostic to failure definitions. HIP-LLM can accommodate any well-defined notion of failure, enabling users to instantiate the model with the criteria most relevant to their domain or evaluation objectives.
How can OPs be specified in practice
OPs are fundamental to software reliability assessment, and the same holds true for LLM reliability. While the hierarchical structure of HIP-LLM provides “interfaces” for integrating OPs, readers may nevertheless question the validity of our OP (which are based on benchmark datasets rather than real operational data). Indeed, in the absence of access to real operational data, this study relies on simulated operational data derived from existing benchmark datasets, treating these datasets as sampling frames that approximate the true unknown operational data distribution. Specifically, we assume task-level distributions within each subdomain and assign operational weights across subdomains and domains. We note, as a reliability assessment modeling paper, accurately estimating OPs from real usage data is beyond our current scope. However, this is a well-recognized challenge in software reliability engineering, for which numerous estimation methods have been proposed (e.g., [25, 26, 28]). The concept of OP has also been discussed and measured in AI and ML contexts [24, 76, 78, 79].
The challenge of modeling OPs does not undermine the framework itself but highlight why the OP is essential: it provides the mechanism to reweight and reinterpret benchmark results in terms of real usage. As more representative datasets or user-usage logs [27] become available, the OP can be updated accordingly, making the reliability estimates more faithful to actual field performance.
We acknowledge that our present treatment of OP uncertainty remains simplified: assuming fixed task distributions in each subdomain and constant (sub-)domain weights. More complex modeling on OP uncertainties [72, 73, 26, 74, 75], e.g., dynamic OPs , will form our important future works.
How to embed prior knowledge
HIP-LLM models uncertainty at three levels: the subdomain success rate ($\theta_{ij}$), how much domains share information ($\mu_{i},\nu_{i}$), and the top-level parameters ($a_{i},b_{i},c_{i},d_{i}$) that control the importance of those priors.
As a Bayesian framework, HIP-LLM requires priors to initiate inference, and these priors may substantially influence the resulting posteriors, particularly when data are limited [80, 81]. The challenge of eliciting meaningful priors is not unique to HIP-LLM but is common to all Bayesian models. Unlike HiBayES [31], which employs noninformative priors, HIP-LLM follows the principles of robust Bayesian analysis [59, 60, 82] by adopting the Imprecise Probability framework. This approach enables the representation of uncertainty about prior beliefs through sets of priors rather than committing to a single distribution. Prior research has shown that Imprecise Probability can facilitate expert elicitation by supporting structured “thought experiments” [65, 83, 84] allowing experts to express vague, partial, or interval-valued prior beliefs. Similarly, HIP-LLM’s prior layer can be informed by such thought experiments and expert judgment. While we acknowledge that eliciting priors in practice remains challenging, HIP-LLM at least offers a practical “what-if” analysis tool by allowing assessors to explore how different prior assumptions might influence posterior reliability claims, for example: if my prior belief is A, then after observing evidence B, the posterior reliability would be C.
6 Conclusion
This paper presents HIP-LLM, a hierarchical Bayesian framework with imprecise probability for evaluating the reliability of LLMs. While HIP-LLM acknowledges the general-purpose nature of contemporary LLMs, it also adheres to the standardized definition of software reliability which inherently depends on specific applications and operational contexts. We reconcile these seemingly conflicting aspects by introducing a multi-level hierarchical structure of evaluation tasks that explicitly models dependencies among subdomains, integrates information through partial pooling, and incorporates OPs at multiple levels of abstraction. Moreover, as a Bayesian inference method, HIP-LLM respects the difficulty of selecting priors. Instead of simply using non-informative priors, HIP-LLM provides the mechanism to embed imprecise prior knowledge and reports posterior envelopes (rather than single distributions). Our experiments on common LLMs demonstrate that these features enable a more nuanced and standardized reliability estimation compared to existing benchmarks and state-of-the-art methods.
While we believe HIP-LLM represents an important first step toward a principled reliability assessment framework for LLMs (and more broadly, for multimodal generative AI) several key limitations remain (cf. Section 5). These include the challenge of formally defining what constitutes a failure for an LLM, accurately estimating OPs, and systematically eliciting imprecise prior knowledge. Future extensions of this work should aim to address these limitations, for example by developing models for uncertain and dynamically evolving OPs and by integrating more refined mechanisms for prior elicitation.
Acknowledgment
SK and XZ have received funding from the European Union’s EU Framework Program for Research and Innovation Europe Horizon (grant agreement No 101202457). XZ’s contribution is also supported by the UK EPSRC New Investigator Award [EP/Z536568/1]. SK’s contribution is supported by the UKRI Future Leaders Fellowship Grant [MR/S035176/1].
Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or European Research Executive Agency (REA). Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Mathematical Derivations
This appendix contains the proofs and supplementary information for the theorems presented in the main text.
A.1 Common Framework and Notation
In a LLM, for a domain $D_{i}$ with subdomains $S_{i1},\dots,S_{in_{i}}$. For each subdomain $j=1,\dots,n_{i}$:
  • 1.
    Parameter: $\boldsymbol{\theta}_{i}=(\theta_{i1},\ldots,\theta_{in_{i}})\in(0,1)^{n_{i}}$ (subdomain reliability)
  • 2.
    Parameter: $n_{F}$ = number of trials in the future
  • 3.
    Data: $C_{i}=\{(C_{ik},N_{ik})\}_{k=1}^{n_{i}}$ (observed correct responses and trial counts)
  • 4.
    Hierarchical prior.
    $C_{ij}\mid\theta_{ij},N_{ij}$ $\sim\text{Binomial}(N_{ij},\theta_{ij}),$
    $\theta_{ij}\mid\mu_{i},\nu_{i}$ $\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Beta}(\mu_{i}\nu_{i},\,(1-\mu_{i})\nu_{i}),$
    $\mu_{i}$ $\sim\text{Beta}(a_{i},b_{i}),\qquad\nu_{i}\sim\text{Gamma}(c_{i},\text{rate}=d_{i}),$
    with $\mu_{i}\in(0,1)$, $\nu_{i}>0$ (domain-level parameters), and hyper- hyperparameters $h_{i}=(a_{i},b_{i},c_{i},d_{i})$.
  • 5.
    Admissible hyperparameters (imprecise prior).
    $\mathcal{A}_{i}=[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}].$
  • 6.
    Common Likelihood:
    $L(\boldsymbol{\theta}_{i})=Pr(C_{i}\mid\boldsymbol{\theta}_{i})=\prod_{j=1}^{n_{i}}\binom{N_{ij}}{C_{ij}}\theta_{ij}^{C_{ij}}(1-\theta_{ij})^{N_{ij}-C_{ij}}$
A.2 Proof of Theorem 1
Theorem 1 For subdomain $S_{ij}$ in domain $D_{i}$, let $C_{i}=\{(C_{ik},N_{ik})\}_{k=1}^{n_{i}}$ be the observed data. Let the admissible set of hyperparameters be
$\mathcal{A}_{i}\;=\;[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}],$
and write $h_{i}=(a_{i},b_{i},c_{i},d_{i})$. Then, for any $h_{i}\in\mathcal{A}_{i}$, the marginal posterior density of $\theta_{ij}$ is
$Pr(\theta_{ij}\mid C_{i},h_{i})=\frac{f_{\mathrm{marg}}(\theta_{ij},C_{i};h_{i})}{Z_{\mathrm{marg}}(h_{i})},$
where $f_{\mathrm{marg}}$ (unnormalized posterior) and $Z_{\mathrm{marg}}$ (normalizing constant) are
$\hskip 28.45274ptf_{\mathrm{marg}}(\theta_{ij},C_{i};h_{i})=$
$\hskip 56.9055pt\int_{0}^{1}\int_{0}^{\infty}\Bigg[\prod_{k\neq j}\int_{0}^{1}d\theta_{ik}\Bigg]L(\boldsymbol{\theta}_{i})\,Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i})\,Pr(\mu_{i},\nu_{i}\mid h_{i})\,d\mu_{i}\,d\nu_{i}$
$Z_{\mathrm{marg}}(h_{i})=\int_{0}^{1}\int_{0}^{\infty}Pr(C_{i}\mid\mu_{i},\nu_{i})\,Pr(\mu_{i},\nu_{i}\mid h_{i})\,d\mu_{i}\,d\nu_{i},$
with $L(\boldsymbol{\theta}_{i})=Pr(C_{i}\mid\boldsymbol{\theta}_{i})$.
The imprecise marginal posterior is characterized by the lower/upper envelopes
$\underline{Pr}(\theta_{ij}\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}Pr(\theta_{ij}\mid C_{i},h_{i}),\qquad\overline{Pr}(\theta_{ij}\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}Pr(\theta_{ij}\mid C_{i},h_{i}).$
Proof of Theorm 1.
We proceed step by step.
Step 1 — Likelihood and hierarchical prior densities. The joint data likelihood factorizes over subdomains:
$L(\boldsymbol{\theta_{i}})=Pr(C_{i}\mid\theta_{i})=\prod_{k=1}^{n_{i}}\binom{N_{ik}}{C_{ik}}\theta_{ik}^{C_{ik}}(1-\theta_{ik})^{N_{ik}-C_{ik}}.$ (5)
Given $(\mu_{i},\nu_{i})$, the prior on $\theta_{i}$ factorizes as
$Pr(\boldsymbol{\theta_{i}}\mid\mu_{i},\nu_{i})=\prod_{k=1}^{n_{i}}\frac{\Gamma(\nu_{i})}{\Gamma(\mu_{i}\nu_{i})\Gamma((1-\mu_{i})\nu_{i})}\,\theta_{ik}^{\mu_{i}\nu_{i}-1}(1-\theta_{ik})^{(1-\mu_{i})\nu_{i}-1}.$
Given the domain-level parameters $(\mu_{i},\nu_{i})$, the subdomain reliabilities are conditionally independent and identically distributed:
$\theta_{ij}\mid\mu_{i},\nu_{i}\;\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\;\mathrm{Beta}\!\big(\mu_{i}\nu_{i},\,(1-\mu_{i})\nu_{i}\big),\qquad j=1,\dots,n_{i}.$
Here $\mu_{i}\in(0,1)$ is the prior mean and $\nu_{i}>0$ is the prior strength (concentration).
For each $k$, the Beta pdf (Single-coordinate density) is:
$Pr(\theta_{ik}\mid\mu_{i},\nu_{i})=\frac{\Gamma(\nu_{i})}{\Gamma(\mu_{i}\nu_{i})\,\Gamma((1-\mu_{i})\nu_{i})}\,\theta_{ik}^{\,\mu_{i}\nu_{i}-1}\,(1-\theta_{ik})^{\,(1-\mu_{i})\nu_{i}-1},\qquad\theta_{ik}\in(0,1).$
Since $\{\theta_{ik}\}_{k=1}^{n_{i}}$ are conditionally independent given $(\mu_{i},\nu_{i})$,
$Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i})$ $=\prod_{k=1}^{n_{i}}Pr(\theta_{ik}\mid\mu_{i},\nu_{i})$ (6)
$=\prod_{k=1}^{n_{i}}\mathrm{Beta}\!\left(\theta_{ik}\mid\mu_{i}\nu_{i},(1-\mu_{i})\nu_{i}\right)$
$=\left(\frac{\Gamma(\nu_{i})}{\Gamma(\mu_{i}\nu_{i})\,\Gamma((1-\mu_{i})\nu_{i})}\right)^{\!n_{i}}\prod_{k=1}^{n_{i}}\theta_{ik}^{\,\mu_{i}\nu_{i}-1}\,(1-\theta_{ik})^{\,(1-\mu_{i})\nu_{i}-1}$
Given hyperparameters $h_{i}=(a_{i},b_{i},c_{i},d_{i})$, we take $\mu_{i}$ and $\nu_{i}$ to be independent:
$Pr(\mu_{i},\nu_{i}\mid h_{i})\;=\;Pr(\mu_{i}\mid h_{i})\,Pr(\nu_{i}\mid h_{i}).$
We assume:
$\mu_{i}\sim\mathrm{Beta}(a_{i},b_{i}),\quad\mu_{i}\in(0,1),\qquad\nu_{i}\sim\mathrm{Gamma}(c_{i},\text{rate}=d_{i}),\quad\nu_{i}>0.$
Explicitly,
$Pr(\mu_{i}\mid h_{i})=\frac{\Gamma(a_{i}+b_{i})}{\Gamma(a_{i})\Gamma(b_{i})}\,\mu_{i}^{\,a_{i}-1}(1-\mu_{i})^{\,b_{i}-1},\qquad Pr(\nu_{i}\mid h_{i})=\frac{d_{i}^{\,c_{i}}}{\Gamma(c_{i})}\,\nu_{i}^{\,c_{i}-1}\,e^{-d_{i}\nu_{i}}$
Thus,
$Pr(\mu_{i},\nu_{i}\mid h_{i})$ $=\mathrm{Beta}(\mu_{i}\mid a_{i},b_{i})\;\mathrm{Gamma}(\nu_{i}\mid c_{i},\text{rate}=d_{i})$ (7)
$=\frac{\Gamma(a_{i}+b_{i})}{\Gamma(a_{i})\,\Gamma(b_{i})}\;\frac{d_{i}^{\,c_{i}}}{\Gamma(c_{i})}\;\mu_{i}^{\,a_{i}-1}(1-\mu_{i})^{\,b_{i}-1}\;\nu_{i}^{\,c_{i}-1}e^{-d_{i}\nu_{i}}$
$=\frac{1}{\mathrm{B}(a_{i},b_{i})}\;\frac{d_{i}^{\,c_{i}}}{\Gamma(c_{i})}\;\mu_{i}^{\,a_{i}-1}(1-\mu_{i})^{\,b_{i}-1}\;\nu_{i}^{\,c_{i}-1}e^{-d_{i}\nu_{i}}$
where $\mathrm{B}(a_{i},b_{i})=\dfrac{\Gamma(a_{i})\Gamma(b_{i})}{\Gamma(a_{i}+b_{i})}$.
The hyperprior factorizes:
$Pr(\mu_{i},\nu_{i}\mid h_{i})=\frac{\Gamma(a_{i}+b_{i})}{\Gamma(a_{i})\Gamma(b_{i})}\mu_{i}^{a_{i}-1}(1-\mu_{i})^{b_{i}-1}\cdot\frac{d_{i}^{\,c_{i}}}{\Gamma(c_{i})}\nu_{i}^{c_{i}-1}e^{-d_{i}\nu_{i}}.$
Step 2 — Normalizing constant. By definition,
$Z_{\mathrm{marg}}(C_{i},h_{i})=\int_{0}^{1}Pr(\theta_{ij}\mid C_{i},h_{i})\,d\theta_{ij}=\int_{0}^{1}\!\!\int_{0}^{\infty}Pr(C_{i}\mid\mu_{i},\nu_{i})\,Pr(\mu_{i},\nu_{i}\mid h_{i})\,d\mu_{i}\,d\nu_{i},$
where
$Pr(C_{i}\mid\mu_{i},\nu_{i})=\int_{(0,1)^{n_{i}}}L(\theta_{i})\,Pr(\theta_{i}\mid\mu_{i},\nu_{i})\,d\theta_{i}.$
Start from the definition:
$Pr(\theta_{ij}\mid C_{i},h_{i})=\frac{f_{\mathrm{marg}}(\theta_{ij},C_{i};h_{i})}{Z_{\mathrm{marg}}(C_{i},h_{i})}$
where
$f_{\mathrm{marg}}(\theta_{ij},C_{i};h_{i})$ $=$
$\int_{0}^{1}\!\!\int_{0}^{\infty}\left[\prod_{k\neq j}\int_{0}^{1}d\theta_{ik}\right]\,L(\boldsymbol{\theta}_{i})\,Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i})\,Pr(\mu_{i},\nu_{i}\mid h_{i})\,d\mu_{i}\,d\nu_{i}$
The posterior integrates to 1:
$\int_{0}^{1}Pr(\theta_{ij}\mid C_{i},h_{i})\,d\theta_{ij}=1\;\;\Longrightarrow\;\;\frac{1}{Z_{\mathrm{marg}}(C_{i},h_{i})}\int_{0}^{1}f_{\mathrm{marg}}(\theta_{ij},C_{i};h_{i})\,d\theta_{ij}=1$
Hence
$Z_{\mathrm{marg}}(C_{i},h_{i})=\int_{0}^{1}f_{\mathrm{marg}}(\theta_{ij},C_{i};h_{i})\,d\theta_{ij}\,$
By expanding the integral:
$Z_{\mathrm{marg}}(C_{i},h_{i})$ $=\int_{0}^{1}\!\!\int_{0}^{1}\!\!\int_{0}^{\infty}\left[\prod_{k\neq j}\int_{0}^{1}d\theta_{ik}\right]\,L(\boldsymbol{\theta}_{i})\,Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i})\,Pr(\mu_{i},\nu_{i}\mid h_{i})\,d\mu_{i}\,d\nu_{i}\,d\theta_{ij}$
$=\int_{0}^{1}\!\!\int_{0}^{\infty}\left[\int_{(0,1)^{n_{i}}}L(\boldsymbol{\theta}_{i})\,Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i})\,d\boldsymbol{\theta}_{i}\right]\,Pr(\mu_{i},\nu_{i}\mid h_{i})\,d\mu_{i}\,d\nu_{i}.$ (8)
By substituting Eqs. 5, 6, 7 into 8 we have:
$Z_{\mathrm{marg}}(C_{i},h_{i})=$ $\left[\prod_{j=1}^{n_{i}}\binom{N_{ij}}{C_{ij}}\right]\;\frac{d_{i}^{\,c_{i}}}{\mathrm{B}(a_{i},b_{i})\,\Gamma(c_{i})}\int_{0}^{1}\!\!\int_{0}^{\infty}\mu_{i}^{\,a_{i}-1}(1-\mu_{i})^{\,b_{i}-1}\;\nu_{i}^{\,c_{i}-1}e^{-d_{i}\nu_{i}}$ (9)
$\hskip 0.0pt\times\left(\frac{\Gamma(\nu_{i})}{\Gamma(\alpha_{i})\,\Gamma(\beta_{i})}\right)^{\!n_{i}}\prod_{j=1}^{n_{i}}\frac{\Gamma\!\big(C_{ij}+\alpha_{i}\big)\;\Gamma\!\big(N_{ij}-C_{ij}+\beta_{i}\big)}{\Gamma\!\big(N_{ij}+\nu_{i}\big)}\;d\nu_{i}\,d\mu_{i}$
Step 3 — Imprecise posterior. For any fixed $h_{i}=(a_{i},b_{i},c_{i},d_{i})\in\mathcal{A}_{i}$, Steps 1–2 give the precise (pointwise) posterior density
$Pr(\theta_{ij}\mid C_{i},h_{i})=\frac{f_{\mathrm{marg}}(\theta_{ij},C_{i};h_{i})}{Z_{\mathrm{marg}}(C_{i},h_{i})},$
The hyperparameter vector $h_{i}$ is not known exactly but only to lie within the admissible hyperrectangle
$\mathcal{A}_{i}=[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}].$
Hence the family $\mathcal{F}_{i}=\{Pr(\theta_{ij}\mid C_{i},h_{i}):h_{i}\in\mathcal{A}_{i}\}$ represents all posterior densities consistent with our prior ignorance about $h_{i}$.
Following the theory of imprecise probabilities, we define the lower and upper posterior densities (envelopes) as the pointwise infimum and supremum over this family:
$\underline{Pr}(\theta_{ij}\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}Pr(\theta_{ij}\mid C_{i},h_{i}),\qquad\overline{Pr}(\theta_{ij}\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}Pr(\theta_{ij}\mid C_{i},h_{i}).$
Operationally, these envelopes are obtained by evaluating the closed-form posterior $Pr(\theta_{ij}\mid C_{i},h_{i})$ at the extremal corners of $\mathcal{A}_{i}$ or through numerical optimization if the extrema occur in the interior. The resulting pair $(\underline{Pr},\overline{Pr})$ bounds all admissible precise posteriors and fully characterizes the imprecise posterior belief about $\theta_{ij}$ given the uncertainty in $(a_{i},b_{i},c_{i},d_{i})$.
A.3 Proof of Theorem 2
Theorem 2 For domain $D_{i}$ with local OP weights $\Omega_{ij}$ (where $\sum_{j=1}^{n_{i}}\Omega_{ij}=1$), let $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ be the domain-level non-failure probability. Define the admissible set of hyper-hyper-parameters
$\mathcal{A}_{i}=[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}],$
and write $h_{i}=(a_{i},b_{i},c_{i},d_{i})$.
Then, for any $h_{i}\in\mathcal{A}_{i}$, the posterior distribution of $p_{i}$ is characterized by its cumulative distribution function:
$F_{p_{i}}(t\mid C_{i},h_{i})$ $=\Pr(p_{i}\leq t\mid C_{i},h_{i})$
$\hskip 56.9055pt=\int_{0}^{1}\int_{0}^{\infty}F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}$
where
  • 1.
    $F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})$ is the conditional CDF of $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ given that $\theta_{ij}\mid\mu_{i},\nu_{i},C_{i}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Beta}(C_{ij}+\mu_{i}\nu_{i},N_{ij}-C_{ij}+(1-\mu_{i})\nu_{i})$ for $j=1,\ldots,n_{i}$,
  • 2.
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})$ is the hyper-posterior obtained via Bayes’ rule:
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})=$
    $\hskip 28.45274pt\frac{Pr(C_{i}\mid\mu_{i},\nu_{i})\,\text{Beta}(\mu_{i}\mid a_{i},b_{i})\,\text{Gamma}(\nu_{i}\mid c_{i},\text{rate}=d_{i})}{\int_{0}^{1}\int_{0}^{\infty}Pr(C_{i}\mid\mu,\nu)\,\text{Beta}(\mu\mid a_{i},b_{i})\,\text{Gamma}(\nu\mid c_{i},\text{rate}=d_{i})\,d\mu\,d\nu}$
The imprecise domain posterior is characterized by CDF envelopes:
$\underline{F}_{p_{i}}(t\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{p_{i}}(t\mid C_{i},h_{i}),\qquad\overline{F}_{p_{i}}(t\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{p_{i}}(t\mid C_{i},h_{i}).$
The domain reliability $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ is a weighted sum of subdomain reliabilities, whose joint posterior distribution $Pr(\boldsymbol{\theta}_{i}\mid C_{i},h_{i})$ has no tractable closed form. However, the hierarchical structure of our model provides a natural decomposition: we can express this joint posterior as a mixture over the shared hyperparameters $(\mu_{i},\nu_{i})$, where conditionally on these hyperparameters, the subdomains become independent with Beta posteriors. This decomposition, formalized in Lemma 7, is the key to deriving the domain-level CDF. The proof proceeds by (1) expressing the domain CDF as an integral of the joint posterior over a weighted-sum constraint region, (2) applying the hierarchical decomposition from the lemma, (3) exchanging the order of integration via Fubini’s theorem, and (4) recognizing the inner integral as a conditional CDF, yielding a tractable mixture representation.
Lemma 7 (Hierarchical decomposition of subdomain joint posterior).
Under the hierarchical model with hyperparameters $h_{i}=(a_{i},b_{i},c_{i},d_{i})$, the joint posterior of subdomain reliabilities $\boldsymbol{\theta}_{i}=(\theta_{i1},\ldots,\theta_{in_{i}})$ admits the decomposition
$Pr(\boldsymbol{\theta}_{i}\mid C_{i},h_{i})=\int_{0}^{1}\int_{0}^{\infty}Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i},$ (10)
where:
  1. 1.
    The conditional distribution factorizes into independent Betas:
    $Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i},C_{i})=\prod_{j=1}^{n_{i}}\text{Beta}(\theta_{ij}\mid C_{ij}+\mu_{i}\nu_{i},N_{ij}-C_{ij}+(1-\mu_{i})\nu_{i}).$
  2. 2.
    The hyper-posterior is obtained via Bayes’ rule:
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})$ $=\frac{Pr(C_{i}\mid\mu_{i},\nu_{i})\,Pr(\mu_{i},\nu_{i}\mid h_{i})}{Pr(C_{i}\mid h_{i})},$ (11)
    $Pr(\mu_{i},\nu_{i}\mid h_{i})$ $=\text{Beta}(\mu_{i}\mid a_{i},b_{i})\times\text{Gamma}(\nu_{i}\mid c_{i},\text{rate}=d_{i}),$
    where the marginal likelihood is
    $Pr(C_{i}\mid\mu_{i},\nu_{i})=\prod_{j=1}^{n_{i}}\binom{N_{ij}}{C_{ij}}\frac{B(C_{ij}+\mu_{i}\nu_{i},N_{ij}-C_{ij}+(1-\mu_{i})\nu_{i})}{B(\mu_{i}\nu_{i},(1-\mu_{i})\nu_{i})},$ (12)
    and the evidence is
    $Pr(C_{i}\mid h_{i})=\int_{0}^{1}\int_{0}^{\infty}Pr(C_{i}\mid\mu,\nu)\,\text{Beta}(\mu\mid a_{i},b_{i})\,\text{Gamma}(\nu\mid c_{i},\text{rate}=d_{i})\,d\mu\,d\nu.$ (13)
Proof of Lemma 7.
The decomposition follows directly from the law of total probability and the hierarchical prior structure. By the tower property of conditional expectation applied to densities:
$Pr(\boldsymbol{\theta}_{i}\mid C_{i},h_{i})=$
$\hskip 28.45274pt\int_{0}^{1}\int_{0}^{\infty}Pr(\boldsymbol{\theta}_{i},\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}=$
$\hskip 28.45274pt\int_{0}^{1}\int_{0}^{\infty}Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}$
The conditional independence of $\{\theta_{ij}\}$ given $(\mu_{i},\nu_{i},C_{i})$ follows from the hierarchical prior:
$\theta_{ij}\mid\mu_{i},\nu_{i}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\text{Beta}(\mu_{i}\nu_{i},(1-\mu_{i})\nu_{i}),\quad j=1,\ldots,n_{i},$
combined with independent Binomial likelihoods $C_{ij}\mid\theta_{ij},N_{ij}\sim\text{Binomial}(N_{ij},\theta_{ij})$. By conjugacy, the Beta-Binomial update yields:
$\theta_{ij}\mid\mu_{i},\nu_{i},C_{i}\sim\text{Beta}(C_{ij}+\mu_{i}\nu_{i},N_{ij}-C_{ij}+(1-\mu_{i})\nu_{i}).$
The hyper-posterior (11) follows from Bayes’ rule. The marginal likelihood (12) is the Beta-Binomial distribution for each subdomain, obtained by integrating the Binomial likelihood against the Beta prior. The evidence (13) is the normalization constant ensuring the hyper-posterior integrates to one. ∎
Proof of Theorem 2.
We derive the posterior CDF of the domain reliability $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ using the hierarchical decomposition from Lemma 7.
Step 1: CDF definition. For any $t\in[0,1]$, the posterior CDF of $p_{i}$ is
$F_{p_{i}}(t\mid C_{i},h_{i})=\Pr(p_{i}\leq t\mid C_{i},h_{i})=\Pr\left(\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}\leq t\,\Big|\,C_{i},h_{i}\right).$
Define the constraint region
$\mathcal{R}_{i}(t):=\left\{\boldsymbol{\theta}_{i}\in(0,1)^{n_{i}}:\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}\leq t\right\}.$
Then
$F_{p_{i}}(t\mid C_{i},h_{i})=\int_{\mathcal{R}_{i}(t)}Pr(\boldsymbol{\theta}_{i}\mid C_{i},h_{i})\,d\boldsymbol{\theta}_{i}.$
Step 2: Apply hierarchical decomposition. Substituting the decomposition from Lemma 7 (equation (10)):
$F_{p_{i}}(t\mid C_{i},h_{i})$ $=\int_{\mathcal{R}_{i}(t)}\left[\int_{0}^{1}\int_{0}^{\infty}Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}\right]d\boldsymbol{\theta}_{i}.$
Step 3: Change of integration order. By Fubini’s theorem (applicable since all integrands are non-negative and integrate to finite values):
$F_{p_{i}}(t\mid C_{i},h_{i})$ $=\int_{0}^{1}\int_{0}^{\infty}\left[\int_{\mathcal{R}_{i}(t)}Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i},C_{i})\,d\boldsymbol{\theta}_{i}\right]Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}.$
Step 4: Conditional CDF. The inner integral is the conditional CDF of $p_{i}$ given $(\mu_{i},\nu_{i})$ and $C_{i}$:
$F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})=\int_{\mathcal{R}_{i}(t)}Pr(\boldsymbol{\theta}_{i}\mid\mu_{i},\nu_{i},C_{i})\,d\boldsymbol{\theta}_{i}=\Pr\left(\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}\leq t\,\Big|\,\mu_{i},\nu_{i},C_{i}\right),$
where $\theta_{ij}\mid\mu_{i},\nu_{i},C_{i}$ are independent Beta random variables as specified in Lemma 7.
Step 5: Mixture representation. Combining Steps 3 and 4:
$F_{p_{i}}(t\mid C_{i},h_{i})=\int_{0}^{1}\int_{0}^{\infty}F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}.$
where
  • 1.
    $F_{p_{i}}(t\mid\mu_{i},\nu_{i},C_{i})$ is the conditional CDF of $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ given that $\theta_{ij}\mid\mu_{i},\nu_{i},C_{i}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Beta}(C_{ij}+\mu_{i}\nu_{i},N_{ij}-C_{ij}+(1-\mu_{i})\nu_{i})$ for $j=1,\ldots,n_{i}$,
  • 2.
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})$ is the hyper-posterior obtained via Bayes’ rule:
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})=$
    $\hskip 28.45274pt\frac{Pr(C_{i}\mid\mu_{i},\nu_{i})\,\text{Beta}(\mu_{i}\mid a_{i},b_{i})\,\text{Gamma}(\nu_{i}\mid c_{i},\text{rate}=d_{i})}{\int_{0}^{1}\int_{0}^{\infty}Pr(C_{i}\mid\mu,\nu)\,\text{Beta}(\mu\mid a_{i},b_{i})\,\text{Gamma}(\nu\mid c_{i},\text{rate}=d_{i})\,d\mu\,d\nu}$
This expresses the domain posterior CDF as a weighted average of conditional CDFs, where the weights are given by the hyper-posterior $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})$.
Step 6: Imprecise probability bounds. For each $h_{i}\in\mathcal{A}_{i}$, the mixture formula defines a valid CDF $F_{p_{i}}(\cdot\mid C_{i},h_{i})$. The imprecise posterior is characterized by pointwise envelopes over the admissible set:
$\underline{F}_{p_{i}}(t\mid C_{i})=\inf_{h_{i}\in\mathcal{A}_{i}}F_{p_{i}}(t\mid C_{i},h_{i}),\quad\overline{F}_{p_{i}}(t\mid C_{i})=\sup_{h_{i}\in\mathcal{A}_{i}}F_{p_{i}}(t\mid C_{i},h_{i}),\quad t\in[0,1]$
A.4 Proof of Theorem 3
Theorem 3 For the LLM system with domain weights $W_{i}$ (where $\sum_{i=1}^{m}W_{i}=1$), let $p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$ be the LLM-level failure probability and $\text{data}=\{C_{1},\ldots,C_{m}\}$ the observed data across all domains. Assume cross-domain independence.
Define the domain-level admissible sets
$\mathcal{A}_{i}=[a_{i}^{\min},a_{i}^{\max}]\times[b_{i}^{\min},b_{i}^{\max}]\times[c_{i}^{\min},c_{i}^{\max}]\times[d_{i}^{\min},d_{i}^{\max}]$
and write $h_{i}=(a_{i},b_{i},c_{i},d_{i})$ for $i=1,\ldots,m$. Define the LLM-level admissible set as the Cartesian product
$\mathcal{A}_{\text{LLM}}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{m},$
and collect the domain hyperparameters as $\mathcal{H}=(h_{1},\ldots,h_{m})\in\mathcal{A}_{\text{LLM}}$.
Then, for any $\mathcal{H}\in\mathcal{A}_{\text{LLM}}$, the posterior distribution of $p_{L}$ is characterized by its cumulative distribution function:
$F_{p_{L}}(t\mid\text{data},\mathcal{H})=Pr(p_{L}\leq t\mid\text{data},\mathcal{H})=$
$\hskip 42.67912pt\int\cdots\int G\big(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data}\big)\prod_{i=1}^{m}Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\prod_{i=1}^{m}d\mu_{i}\,d\nu_{i}$
where
  • 1.
    $G(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data})$ is the conditional CDF of $p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$ given all hyperparameters, defined as
    $G\big(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data}\big)=\int_{\mathcal{R}_{L}(t)}\prod_{i=1}^{m}f_{p_{i}}(p_{i}\mid\mu_{i},\nu_{i},C_{i})\,dp_{1}\cdots dp_{m},$
    where $\mathcal{R}_{L}(t):=\{(p_{1},\ldots,p_{m})\in(0,1)^{m}:\sum_{i=1}^{m}W_{i}p_{i}\leq t\}$, and $f_{p_{i}}(\cdot\mid\mu_{i},\nu_{i},C_{i})$ is the conditional density of $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ under $\theta_{ij}\mid\mu_{i},\nu_{i},C_{i}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Beta}(C_{ij}+\mu_{i}\nu_{i},N_{ij}-C_{ij}+(1-\mu_{i})\nu_{i})$,
  • 2.
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})$ is the domain-level hyper-posterior for domain $i$:
    $Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})=$
    $\hskip 28.45274pt\frac{Pr(C_{i}\mid\mu_{i},\nu_{i})\,\text{Beta}(\mu_{i}\mid a_{i},b_{i})\,\text{Gamma}(\nu_{i}\mid c_{i},\text{rate}=d_{i})}{\int_{0}^{1}\int_{0}^{\infty}Pr(C_{i}\mid\mu,\nu)\,\text{Beta}(\mu\mid a_{i},b_{i})\,\text{Gamma}(\nu\mid c_{i},\text{rate}=d_{i})\,d\mu\,d\nu}$
  • 3.
    Cross-domain independence ensures $Pr(\{\mu_{i},\nu_{i}\}_{i=1}^{m}\mid\text{data},\mathcal{H})=\prod_{i=1}^{m}Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})$.
The imprecise LLM posterior is characterized by CDF envelopes:
$\underline{F}_{p_{L}}(t\mid\text{data})=\inf_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{p_{L}}(t\mid\text{data},\mathcal{H})$
$\overline{F}_{p_{L}}(t\mid\text{data})=\sup_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{p_{L}}(t\mid\text{data},\mathcal{H})$
Proof of Theorem 3.
We derive the posterior distribution of the LLM reliability $p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$ by characterizing its cumulative distribution function, building on the domain-level results from Theorem 2.
Step 1: CDF definition. For any $t\in[0,1]$, the posterior CDF of $p_{L}$ is
$F_{p_{L}}(t\mid\text{data},\mathcal{H})=\Pr(p_{L}\leq t\mid\text{data},\mathcal{H})=\Pr\left(\sum_{i=1}^{m}W_{i}p_{i}\leq t\,\Big|\,\text{data},\mathcal{H}\right).$
Define the region
$\mathcal{R}_{L}(t):=\left\{(p_{1},\ldots,p_{m})\in(0,1)^{m}:\sum_{i=1}^{m}W_{i}p_{i}\leq t\right\}.$
Then
$F_{p_{L}}(t\mid\text{data},\mathcal{H})=\int_{\mathcal{R}_{L}(t)}\prod_{i=1}^{m}Pr(p_{i}\mid C_{i},h_{i})\,dp_{1}\cdots dp_{m},$
where cross-domain independence ensures the joint density factorizes as
$Pr(p_{1},\ldots,p_{m}\mid\text{data},\mathcal{H})=\prod_{i=1}^{m}Pr(p_{i}\mid C_{i},h_{i}).$
Step 2: Domain-level mixture representation. From Theorem 2, each domain posterior can be written as
$Pr(p_{i}\mid C_{i},h_{i})=\int_{0}^{1}\int_{0}^{\infty}f_{p_{i}}(p_{i}\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i},$
where $f_{p_{i}}(\cdot\mid\mu_{i},\nu_{i},C_{i})$ is the conditional density of $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$ given that
$\theta_{ij}\mid\mu_{i},\nu_{i},C_{i}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Beta}(C_{ij}+\mu_{i}\nu_{i},N_{ij}-C_{ij}+(1-\mu_{i})\nu_{i}),\quad j=1,\ldots,n_{i}.$
Step 3: Hierarchical decomposition across domains. Substituting the domain-level mixture into the LLM-level CDF:
$F_{p_{L}}(t\mid\text{data},\mathcal{H})=$
$\hskip 28.45274pt\int_{\mathcal{R}_{L}(t)}\prod_{i=1}^{m}\left[\int_{0}^{1}\int_{0}^{\infty}f_{p_{i}}(p_{i}\mid\mu_{i},\nu_{i},C_{i})\,Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\,d\mu_{i}\,d\nu_{i}\right]dp_{1}\cdots dp_{m}$
Step 4: Change of integration order. By Fubini’s theorem (applicable because all integrands are non-negative and integrate to finite values), we can exchange the order of integration:
$F_{p_{L}}(t\mid\text{data},\mathcal{H})$ $=\int_{0}^{1}\cdots\int_{0}^{1}\int_{0}^{\infty}\cdots\int_{0}^{\infty}\left[\int_{\mathcal{R}_{L}(t)}\prod_{i=1}^{m}f_{p_{i}}(p_{i}\mid\mu_{i},\nu_{i},C_{i})\,dp_{1}\cdots dp_{m}\right]$
$\qquad\qquad\qquad\qquad\qquad\qquad\times\prod_{i=1}^{m}Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\prod_{i=1}^{m}d\mu_{i}\,d\nu_{i}.$
Step 5: Conditional CDF. The inner integral is the conditional CDF of $p_{L}$ given all hyperparameters:
$G\big(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data}\big)$ $:=\int_{\mathcal{R}_{L}(t)}\prod_{i=1}^{m}f_{p_{i}}(p_{i}\mid\mu_{i},\nu_{i},C_{i})\,dp_{1}\cdots dp_{m}$
$=\Pr\left(\sum_{i=1}^{m}W_{i}p_{i}\leq t\,\Big|\,\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data}\right),$
where $p_{i}\mid\mu_{i},\nu_{i},C_{i}$ are independent across domains, each distributed according to the weighted sum of independent Betas as specified in Step 2.
Step 6: Mixture representation. Combining Steps 4 and 5:
$F_{p_{L}}(t\mid\text{data},\mathcal{H})=$
$\hskip 28.45274pt\int\cdots\int G\big(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data}\big)\prod_{i=1}^{m}Pr(\mu_{i},\nu_{i}\mid C_{i},h_{i})\prod_{i=1}^{m}d\mu_{i}\,d\nu_{i}$ (14)
where the integrals range over $\mu_{i}\in(0,1)$ and $\nu_{i}\in(0,\infty)$ for all $i=1,\ldots,m$.
This expresses the LLM posterior CDF as a weighted average of conditional CDFs, where the weights are given by the product of independent domain-level hyper-posteriors.
Step 7: Imprecise probability bounds. For each $\mathcal{H}\in\mathcal{A}_{\text{LLM}}$, the formula above defines a valid CDF $F_{p_{L}}(\cdot\mid\text{data},\mathcal{H})$. The imprecise posterior is characterized by pointwise envelopes:
$\underline{F}_{p_{L}}(t\mid\text{data})=\inf_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{p_{L}}(t\mid\text{data},\mathcal{H}),$
$\overline{F}_{p_{L}}(t\mid\text{data})=\sup_{\mathcal{H}\in\mathcal{A}_{\text{LLM}}}F_{p_{L}}(t\mid\text{data},\mathcal{H})$
where, $t\in[0,1]$
A.5 Reliability for Multiple Consecutive Operations
Theorems 1, 2, and 3 establish posterior distributions for the reliability parameters $\theta_{ij}$, $p_{i}$, and $p_{L}$ at the subdomain, domain, and LLM levels, respectively. These parameters represent the probability of success on a single future operation. However, in practical reliability assessment, we often need to evaluate performance over multiple consecutive operations.
This appendix clarifies the mathematical framework for extending single-operation reliability to $n_{F}$ consecutive operations, which is formalized in Theorems 4, 5, and 6. We explain:
  1. 1.
    Why reliability over $n_{F}$ operations takes the form $\theta_{ij}^{n_{F}}$ (or $p_{i}^{n_{F}}$, $p_{L}^{n_{F}}$)
  2. 2.
    Why we characterize these distributions through CDFs rather than closed-form densities
  3. 3.
    How to derive the CDF integral formulas
  4. 4.
    How this framework extends across all hierarchical levels
A.5.1 Definition: $n_{F}$-Operation Reliability
At the subdomain level, we define reliability for $n_{F}$ consecutive future operations as:
$R_{ij}(n_{F})=\theta_{ij}^{n_{F}}$
Assume that:
  • 1.
    Each operation in subdomain $S_{ij}$ succeeds independently with probability $\theta_{ij}$
  • 2.
    Operations are identically distributed (i.i.d. assumption)
  • 3.
    We observe $n_{F}$ consecutive operations
Under these assumptions, the probability that all $n_{F}$ operations succeed is:
$\Pr(\text{all $n_{F}$ operations succeed})=\underbrace{\theta_{ij}\times\theta_{ij}\times\cdots\times\theta_{ij}}_{n_{F}\text{ times}}=\theta_{ij}^{n_{F}}$
Because $\theta_{ij}$ is uncertain—it is a random variable with posterior distribution $Pr(\theta_{ij}\mid C_{i},h_{i})$ from Theorem 1—the $n_{F}$-operation reliability $R_{ij}(n_{F})=\theta_{ij}^{n_{F}}$ is also a random variable. We must therefore characterize its full posterior distribution.
A.5.2 Why the CDF Approach
Unlike the posteriors in Theorems 13, which characterize reliability for a single future operation, the distribution of $R_{ij}(n_{F})=\theta_{ij}^{n_{F}}$ generally does not have a closed-form probability density function, even when $\theta_{ij}$ follows a Beta distribution. This is because:
  • 1.
    The transformation $\theta_{ij}\mapsto\theta_{ij}^{n_{F}}$ is nonlinear for $n_{F}\neq 1$
  • 2.
    At higher hierarchical levels (domain, LLM), we have weighted sums of dependent random variables raised to powers, making closed forms even less tractable
Therefore, we characterize the distribution of $R_{ij}(n_{F})$ through its cumulative distribution function (CDF), which has a tractable integral representation and is sufficient for all practical reliability calculations (e.g., computing probabilities, quantiles, expected values).
A.5.3 Deriving the CDF Formula
What does CDF answers: “What is the probability that $n_{F}$-operation reliability is at most rr?”
$F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i})=\Pr(R_{ij}(n_{F})\leq r\mid C_{i},h_{i})=\Pr(\theta_{ij}^{n_{F}}\leq r\mid C_{i},h_{i})$
To compute this probability, we use the following key observation. If $\theta_{ij}^{n_{F}}\leq r$, then taking the $n_{F}$-th root of both sides:
$\theta_{ij}\leq r^{1/n_{F}}$
Since $\theta_{ij}\in(0,1)$ and $r\in[0,1]$, both sides are positive, and the function $x\mapsto x^{1/n_{F}}$ is monotonically increasing on $[0,1]$ for $n_{F}>0$. Therefore, the inequality is preserved under this transformation.
It follows that:
$\Pr(\theta_{ij}^{n_{F}}\leq r\mid C_{i},h_{i})=\Pr(\theta_{ij}\leq r^{1/n_{F}}\mid C_{i},h_{i})$
The right-hand side is simply the CDF of the posterior distribution $Pr(\theta_{ij}\mid C_{i},h_{i})$ from Theorem 1, evaluated at $r^{1/n_{F}}$:
$\Pr(\theta_{ij}\leq r^{1/n_{F}}\mid C_{i},h_{i})=\int_{0}^{r^{1/n_{F}}}Pr(\theta_{ij}\mid C_{i},h_{i})\,d\theta_{ij}$
This gives us the formula in Theorem 4.
A.5.4 Extension to Domain and LLM Levels
The same logic extends to domain and LLM levels:
  • 1.
    Domain level: $R_{i}(n_{F})=p_{i}^{n_{F}}$ where $p_{i}=\sum_{j=1}^{n_{i}}\Omega_{ij}\theta_{ij}$
  • 2.
    LLM level: $R_{L}(n_{F})=p_{L}^{n_{F}}$ where $p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$
At these levels, closed forms are even less available because we must:
  1. 1.
    Compute the distribution of weighted sums of dependent Beta random variables (from Theorems 2 and 3)
  2. 2.
    Apply the power transformation to obtain $R_{i}(n_{F})$ or $R_{L}(n_{F})$
Both steps lack analytical solutions, so we compute CDFs via numerical integration or Monte Carlo sampling.
Appendix B Global numerical settings used in all figures
All figures in this paper are produced under the same imprecise hierarchical Bayes setup and numerical settings:
  • 1.
    Domains/subdomains. D1 (Coding) = {MBPP, DS-1000},  D2 (Reasoning) = {BoolQ, RACE-H}.
  • 2.
    Aggregation weights. Within-domain weights $\Omega_{1}=[0.204,\,0.796]$ (MBPP, DS-1000) and $\Omega_{2}=[0.483,\,0.517]$ (BoolQ, RACE-H); LLM-level weights W=[0.149, 0.851]W=[0.149,\,0.851] (D1, D2).
  • 3.
    Hyperpriors and imprecision ranges (per domain $i$).
    $\mu_{i}\sim\mathrm{Beta}(a_{i},b_{i})$
    ,  $\nu_{i}\sim\mathrm{Gamma}(c_{i},\text{rate}=d_{i})$,  with $a_{i},b_{i}\in[1,12]$ and $c_{i},d_{i}\in[1,25]$.
  • 4.
    simulation For each domain and each sampled hyperparameter configuration: (1) sample $(\mu_{i},\nu_{i})$ from the discretized posterior; (2) draw subdomain accuracies $\theta_{ij}$ from their Beta posteriors; (3) aggregate to $p_{i}$ and $p_{L}$; (4) compute either empirical CDFs of $Z\in\{\theta_{ij},p_{i},p_{L}\}$ on a fixed grid, or Monte Carlo expectations such as $\widehat{\mathbb{E}}[R_{L}(n_{F})]=\frac{1}{S}\sum_{s=1}^{S}\big(p_{L}^{(s)}\big)^{n_{F}}$. Repeating over $K$ hyperparameter configurations yields a family of CDFs/expectations; the plotted bands are the pointwise min–max across configurations (epistemic uncertainty).
  • 5.
    CDF envelopes. For any quantity $Z\in\{\theta_{ij},p_{i},p_{L}\}$, we compute an empirical CDF for each configuration, and plot the pointwise min–max envelope $F_{\min}(t),F_{\max}(t)]$ across configurations.
Appendix C List of Mathematical Notations
Symbol Description
$\mathcal{X}$ Input-space of all possible tasks for an LLM
$\pi$ Operational Profile (OP), probability distribution over tasks
$n$ Number of future tasks
$n_{F}$ Number of consecutive future operations
$I(x_{t})$ Indicator function: 1 for success, 0 for failure on task $t$
$R(n,\pi)$ LLM reliability: probability of failure-free operation over $n$ tasks
$M$ Number of LLM models being evaluated
$m$ Number of independent domains
$D_{i}$ Domain $i$ ($i=1,\ldots,m$)
$n_{i}$ Number of subdomains in domain $i$
$S_{ij}$ Subdomain $j$ in domain $i$ ($j=1,\ldots,n_{i}$)
$C_{ij}$ Number of correct responses in subdomain $S_{ij}$
$N_{ij}$ Number of trials (tasks) in subdomain $S_{ij}$
$C_{i}$ Set of observed data in domain $i$: $\{(C_{ik},N_{ik})\}_{k=1}^{n_{i}}$
data Observed data across all domains: $\{C_{1},\ldots,C_{m}\}$
$\theta_{ij}$ Subdomain reliability (success probability) for $S_{ij}$
$\boldsymbol{\theta}_{i}$ Vector of subdomain reliabilities: $(\theta_{i1},\ldots,\theta_{in_{i}})$
$p_{i}$ Domain-level reliability
$p_{L}$ LLM-level reliability
$R_{ij}(n_{F})$ Subdomain reliability for $n_{F}$ consecutive operations
$R_{i}(n_{F})$ Domain reliability for $n_{F}$ consecutive operations
$R_{L}(n_{F})$ LLM reliability for $n_{F}$ consecutive operations
$\Omega_{ij}$ Operational weight for subdomain $j$ in domain $i$
$W_{i}$ Operational weight for domain $i$
$\mu_{i}$ Expected reliability (prior mean) for domain $i$
$\nu_{i}$ Prior strength (concentration parameter) for domain $i$
$a_{i},b_{i}$ Hyperparameters for Beta prior on $\mu_{i}$
$c_{i},d_{i}$ Hyperparameters for Gamma prior on $\nu_{i}$
$h_{i}$ Tuple of hyperparameters: $(a_{i},b_{i},c_{i},d_{i})$
$\mathcal{A}_{i}$ Admissible set of hyperparameters for domain $i$
$\mathcal{H}$ Tuple of all domain hyperparameters: $(h_{1},\ldots,h_{m})$
$\mathcal{A}_{\text{LLM}}$ Admissible set at LLM level: $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{m}$
$F_{p_{i}}(t\mid C_{i},h_{i})$ CDF of domain reliability $p_{i}$
$\underline{F}_{p_{i}}(t\mid C_{i})$ Lower CDF envelope for domain reliability
$\overline{F}_{p_{i}}(t\mid C_{i})$ Upper CDF envelope for domain reliability
$F_{p_{L}}(t\mid\text{data},\mathcal{H})$ CDF of LLM reliability $p_{L}$
$\underline{F}_{p_{L}}(t\mid\text{data})$ Lower CDF envelope for LLM reliability
$\overline{F}_{p_{L}}(t\mid\text{data})$ Upper CDF envelope for LLM reliability
$F_{R_{ij}(n_{F})}(r\mid C_{i},h_{i})$ CDF of subdomain $n_{F}$-operation reliability
$F_{R_{i}(n_{F})}(r\mid C_{i},h_{i})$ CDF of domain $n_{F}$-operation reliability
$F_{R_{L}(n_{F})}(r\mid\text{data},\mathcal{H})$ CDF of LLM $n_{F}$-operation reliability
Symbol Description
$f_{\text{marg}}(\theta_{ij},C_{i};h_{i})$ Unnormalized marginal posterior for $\theta_{ij}$
$Z_{\text{marg}}(h_{i})$ Normalizing constant for marginal posterior
$f_{p_{i}}(\cdot\mid\mu_{i},\nu_{i},C_{i})$ Conditional density of domain reliability
$G(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data})$ Conditional CDF of $p_{L}$ given all hyperparameters
$\mathcal{R}_{i}(t)$ Constraint region for domain $i$
$\mathcal{R}_{L}(t)$ Constraint region for LLM: $\{(p_{1},\ldots,p_{m})$
$f_{\text{marg}}(\theta_{ij},C_{i};h_{i})$ Unnormalized marginal posterior for $\theta_{ij}$
$Z_{\text{marg}}(h_{i})$ Normalizing constant for marginal posterior
$f_{p_{i}}(\cdot\mid\mu_{i},\nu_{i},C_{i})$ Conditional density of domain reliability
$G(t\mid\{\mu_{i},\nu_{i}\}_{i=1}^{m},\text{data})$ Conditional CDF of $p_{L}$ given all hyperparameters
$\mathcal{R}_{i}(t)$ Constraint region for domain $i$
$\mathcal{R}_{L}(t)$ Constraint region for LLM: $\{(p_{1},\ldots,p_{m})$